### v3

# PINN-MEP: Continuous Neural Representations for Minimum-Energy Path Discovery in Molecular Systems

Abstract
Characterizing conformational transitions in physical systems remains a fundamental challenge in the computational sciences. Traditional sampling methods like molecular dynamics (MD) or MCMC often struggle with the high-dimensional nature of molecular systems and the high energy barriers of transitions between stable states. While these transitions are rare events in simulation timescales, they often represent the most biologically significant processes, for example, the conformational change of an ion channel protein from its closed to open state, which controls cellular ion flow and is crucial for neural signaling. Such transitions in real systems may take milliseconds to seconds but could require months or years of continuous simulation to observe even once. We present a method that reformulates transition path generation as a continuous optimization problem solved through physics-informed neural networks (PINNs) inspired by string methods for minimum-energy path (MEP) generation. By representing transition paths as implicit neural functions and leveraging automatic differentiation with differentiable molecular dynamics force fields, our method enables the efficient discovery of physically realistic transition pathways without requiring expensive path sampling. We demonstrate our method’s effectiveness on two proteins, including an explicitly hydrated bovine pancreatic trypsin inhibitor (BPTI) system with over 8,300 atoms. Our approach reproduces the conformational change captured in a landmark millisecond-scale explicit-solvent MD simulation (Shaw et al., 2010), while achieving remarkable computational efficiency, requiring only force field evaluations compared to the approximately 412 billion evaluations in the original study. This represents a nearly six orders of magnitude reduction, allowing us to generate the transition pathway in just 15 minutes on a standard GPU rather than weeks on specialized hardware.
1 Introduction
Understanding molecular conformational changes is fundamental to understanding biological processes, from enzyme catalysis to protein folding. Although molecular dynamics (MD) simulations provide unprecedented atomic-level detail of these processes (Dror et al., 2012), capturing rare transition events remains computationally challenging due to the inherent timescale gap between molecular motion and conformational changes. This difficulty arises from transition rates between stable states being exponentially suppressed by the height of the energy barriers that separate them, as described by transition state theory (Eyring, 1935). As a result, direct simulation methods spend the vast majority of computational effort sampling stable states, making the observation of transitions between them prohibitively expensive.
The search for efficient methods to identify transition pathways between stable states has a rich history in computational physics. Early theoretical work established key principles through the Freidlin-Wentzell (Freidlin & Wentzell, 1998) and Onsager-Machlup (Onsager & Machlup, 1953) functionals, which connect the most probable transition paths to minimizers of specific action functionals in different limiting regimes. Building on these foundations, the minimum-energy path (MEP) between stable molecular states represents the most probable transition pathway in the limit of high friction. This path reveals the mechanism of conformational change, identifies key transition states and intermediate configurations, and can thus be highly informative about the system’s dynamics. Moreover, the MEP can serve as a scaffold for subsequent sampling of the full transition path ensemble, making it a powerful tool for understanding molecular dynamics. Methods such as the string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Petersen et al., 2024; Dellago et al., 1998) approximate these paths through a discrete series of molecular configurations, connected by artificial spring forces to ensure continuity or as splines. Although grounded in solid theoretical principles, these approaches face fundamental challenges stemming from their optimization in Cartesian molecular coordinates. The discrete chain of conformations requires periodic reparameterization to maintain even spacing, leading to a complex alternation between energy minimization and geometric constraints. Additional limitations include the difficulty of generating good initial transition guesses and limited parallelization, which restricts the scale of practical applications. While this formulation has proven useful for many small-scale systems, it has not yet fully benefited from modern optimization approaches that could address these challenges.
Concurrent advances in scientific machine learning and molecular simulation have created opportunities for novel approaches to this challenge. Physics-informed neural networks (PINNs) have revolutionized our ability to solve complex physical problems by embedding differential equations directly into neural architectures (Raissi et al., 2019). Additionally, the development of differentiable molecular force fields (Wang et al., 2023) has enabled seamless integration of physical forcefield models used in MD simulations with modern machine learning frameworks, allowing efficient computation of derivatives needed for path optimization.
The last decade has seen numerous approaches emerge to tackle the transition path sampling problem. Traditional shooting and other sampling methods (Jung et al., 2023; Lazzeri et al., 2023; Bolhuis et al., 2002; Dellago et al., 1998; Laio & Parrinello, 2002; Torrie & Valleau, 1977) rely on expensive molecular dynamics simulations, require initial MD-harvested trajectories to begin with or preexisting knowledge of the transition dynamics, and face inherent tradeoffs between rejection rates and path diversity. More recent machine learning approaches have explored several directions: variational approaches using Doob’s h-transform (Du et al., 2024; Das & Limmer, 2019; Singh & Limmer, 2023), stochastic control formulations (Yan et al., 2022; Holdijk et al., 2023), and even attempts to sample full transition trajectories directly using diffusion models (Petersen et al., 2023; Han et al., 2024; Jing et al., 2024). Notably, Du et al. (2024) use a similar neural representation of transition paths but with a different objective, targeting the complete transition path ensemble via Doob’s h-transform. This approach provides rich statistical information about transition mechanisms. Most closely related to our method, Ramakrishnan et al. (2025) employ a continuous neural parameterization of minimum-energy paths inspired by Nudged Elastic Band (NEB), incorporating specialized components for tangential forces, path velocity, and transition state energetics. Another promising concurrent approach by Raja et al. (2025) creatively leverages pretrained diffusion models as a sampler of initial transition conformations and as a forcefield within a chain-of-states framework to generate plausible transition distributions. While each approach offers valuable insights into neural network-based path discovery, our method focuses on a streamlined optimization scheme that efficiently scales to large biomolecular systems with explicit solvent while maintaining fast training times.
Here, we introduce an approach focused on efficient minimum-energy path discovery as a foundation for understanding molecular transition mechanisms. Our method represents transition paths as continuous, differentiable neural functions that map a progress coordinate to molecular configurations. This representation, inspired by recent advances in neural implicit representations (Mildenhall et al., 2020; Sitzmann et al., 2020), naturally incorporates physical constraints through a physics-informed neural network (PINN) framework. By directly targeting the minimum-energy path through energy minimization, we achieve remarkable computational efficiency, enabling applications to complex biomolecular systems in minutes on standard hardware while extracting the most physically relevant information about transition mechanisms. The resulting pathways can serve as valuable starting points for established transition path sampling methods (Dellago et al., 1998; Bolhuis et al., 2002) when exploration of the complete ensemble is desired.
2 Theory and Methods
2.1 Derivation of the Energy-Based Loss Function and Neural Implementation
To our loss function, we lean on alternative formulations of the Onsager-Machlup action functional as derived by Vanden-Eijnden & Heymann (2008); Olender & Elber (1997) in the context of the string methods. We begin with a molecular system evolving under overdamped Langevin dynamics, a well-established model for molecular motion in the high-friction regime characteristic of biomolecular systems:
represents the potential energy function, is Gaussian white noise modeling thermal fluctuations, and physical constants are absorbed into the nondimensionalization. To derive the path probability, we discretize the time interval into steps of duration , approximating the dynamics as:
Where are independent standard Gaussian random variables. The conditional probability of transitioning from to follows a Gaussian distribution:
To obtain the probability of an entire continuous trajectory, we take the product of these conditional probabilities over all time steps and then take the limit as . This yields the path probability:
The exponent term can be recognized as the Onsager-Machlup action functional, which is defined as:
Thus, maximizing the probability of a transition path (subject to fixed endpoints and ) is equivalent to minimizing this action functional. The action penalizes paths that deviate from following the deterministic drift . To further simplify, we expand the integrand of the action functional:
Applying the Cauchy-Schwarz inequality to the kinetic and potential terms, we can establish an important relationship:
with equality holding if and only if is parallel (or antiparallel) to . When the path is reparameterized so that velocity consistently aligns with the gradient, the action is minimized, yielding the most likely transition pathway. For such pathways where is everywhere proportional to , we can transform the time integral into a purely geometric one. However, to maintain this parallel alignment, the path length itself must be subject to optimization, as a fixed-length path would generally not satisfy the parallelity condition. By defining the arc-length element along the spatial curve traced out by , we obtain:
This geometric reformulation connects directly to the concept of a minimum-energy path (MEP), which follows the potential energy landscape’s gradient while avoiding high-energy regions. The MEP represents the most probable transition mechanism in the overdamped limit, making it an ideal target for our optimization.
We represent the transition path as a continuous neural mapping for such that and , where represents the network parameters. To transform the geometric line integral into a practical loss function, we first discretize the path and apply Taylor expansion to relate gradient magnitudes to energy differences:
where and . Using the triangle inequality, we can establish that minimizing provides an upper bound on this action, which yields our neural network loss function:
A complete derivation with all mathematical details is provided in Appendix A.1. By minimizing , we drive the path representation to discover low-energy regions connecting stable states, approximating the MEP predicted by the Onsager-Machlup framework.
2.1.1 Neural Representation of minimum-energy Paths
Drawing inspiration from the string method formalism, we propose representing the transition path using an implicit neural parameterization, similar to the mean parameterization of Du et al. (Du et al., 2024). While our approaches share similar representations, they differ in objectives and computational efficiency (see Appendix B.1 for theoretical connections and performance comparisons, respectively). Specifically, we construct a continuous function mapping a progress coordinate to molecular configurations , where is the number of atoms:
Here is a neural network with parameters , and are basis functions that satisfy the endpoint constraints , , , and , ensuring that and regardless of the network’s output. While simple linear functions and can be used, other basis functions with appropriate boundary conditions are also effective. The blending function controls the network’s contribution along the path, typically designed to allow maximum influence in the transition region while smoothly vanishing at endpoints (e.g., ).
2.1.2 Path Optimization
Rather than alternating between evolution and reparameterization steps as in the classical string method, we minimize directly through stochastic gradient descent. For each batch of paths:
The implicit neural representation naturally handles reparameterization constraints through its continuous parameterization. Unlike traditional chain-of-states methods, explicit spring-like continuity terms are unnecessary—the inherent continuity of , combined with smooth blending functions, ensures path continuity without additional constraints.
2.2 Established Approaches to Minimum-Energy Path Finding
Classical approaches to minimum-energy path discovery operate directly in Cartesian atomic coordinates, predating differentiable molecular force fields (Wang et al., 2023) that enable our neural network representation. This direct optimization faces significant challenges as steric clashes create extreme energy barriers that are difficult to overcome with local optimization. The string method and chain-of-states approaches discussed below represent broad families with numerous variations, extensively cataloged in reviews such as "Transition-Path Theory and Path-Finding Algorithms for the Study of Rare Events" by E & Vanden-Eijnden (2010). These established techniques illustrate both foundational concepts and limitations that our approach addresses.
2.2.1 String Method
The string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Ramakrishnan et al., 2025) determines MEPs by evolving a curve in configuration space according to physical forces while maintaining proper parameterization. Given the same endpoints and used in our formulation, the method seeks a curve (string) parameterized by that connects these states. The MEP satisfies:
where is the component of the force perpendicular to the unit tangent vector along the string. This orthogonality condition has an intuitive interpretation: at each point along the MEP, the forces acting on the system are entirely parallel to the path itself, with no perpendicular components trying to push the system away from the path. This represents a ridge or valley in the energy landscape—the system naturally follows the path without being forced sideways.
The simplest dynamics for evolving a curve toward this MEP is given by:
Here is a Lagrange multiplier term enforcing the parameterization constraint, typically equal arc-length parameterization where (constant) for all . In practice, the string is discretized into images and evolved through a two-step procedure:
1. Evolution step:
2. Reparameterization step to enforce equal spacing:
This two-step procedure converges to a discretized MEP satisfying each image’s local orthogonality condition . Unlike the neural approach presented in this paper, which handles parameterization implicitly, this method requires explicit reparameterization steps that can introduce computational inefficiencies.
2.2.2 Chain-of-States Methods
An alternative perspective views MEP finding as optimization of a chain of states (Pratt, 1986; Elber & Karplus, 1987; Ulitsky & Elber, 1990; Jónsson et al., 1998; Henkelman et al., 2000; Gillilan & Lilien, 2004; Sheppard et al., 2008). Given a sequence with and , the objective function is defined as:
where and is a spring constant. The spring term enforces continuity of the path while allowing flexibility in the discretization. However, this formulation introduces a well-known limitation: the spring forces tend to cause "corner cutting," where the path takes shortcuts through high-energy regions rather than following the low-energy valleys representing physically meaningful transitions. The neural representation described earlier avoids this limitation through its natural continuity properties.
At convergence, both formulations yield a discretized MEP satisfying at each image, the same condition that PINN-MEP targets through energy minimization, but with the advantage of continuous representation and gradient-based optimization.
3 Experiments
We evaluate our method on two molecular systems: the AIB9 peptide and Bovine Pancreatic Trypsin Inhibitor (BPTI). These systems represent different scales of complexity and serve to validate our method’s ability to discover physically meaningful transition paths. In Appendix A.2, we compare and benchmark our method to the related approach outlined in (Du et al., 2024), and in Appendix B.1, we present comprehensive ablation studies analyzing the impact of different embedding strategies, interpolation methods, network architectures, and regularization approaches.
3.1 AIB9 Transition Path Discovery
We first examine the AIB9 system, a 9-residue artificial protein with 129 atoms that exhibits two well-defined metastable states, making it an ideal test case for transition path methods. This system effectively bridges abstract models and complex proteins: it is small enough for extensive reference simulations yet complex enough to exhibit realistic conformational changes; its energy landscape is well-characterized; and its transitions are easily visualized through dihedral angle projections. Unlike artificial toy problems, AIB9 allows direct validation with actual biomolecular force fields. We use the AMBER ff15ipq-m force field for protein mimetics (Bogetti et al., 2020) implemented in DMFF (Wang et al., 2023), and we generate multiple minimum-energy paths between the system’s two metastable states.
Figure 2 shows the discovered transition paths projected onto the central residue’s - dihedral angle space. The method identifies multiple physically plausible pathways between the states, also observed with long MD simulations. For each path, different conformations of the states were picked as start and end points; this, along with the neural network’s random initialization, leads to the discovery of multiple different transitions.
4 BPTI Conformational Change Pathway
Bovine Pancreatic Trypsin Inhibitor (BPTI) is a clinically significant protein used as an anticoagulant in medical procedures. Beyond its therapeutic applications, BPTI has become a fundamental model system for studying protein dynamics due to its moderate size (58 residues) and complex conformational changes involving disulfide bond rearrangements.
To investigate BPTI’s conformational transitions, we analyze five key states from a landmark molecular dynamics (MD) simulation performed by D.E. Shaw Research on their Anton supercomputer (Shaw et al., 2010). These publicly available snapshots provide rare insight into BPTI’s transition mechanisms. Full details of our BPTI system preparation, explicit solvent shell construction, and hybrid neural network architecture for handling water molecules are provided in Appendix B.2.
We generate a transition MEP connecting the initial and final BPTI states from the D.E. Shaw simulation. The intermediate snapshots from the D.E. Shaw simulation serve as validation points; they were not used as inputs but represent physically realized conformations from the whole trajectory. To analyze this path, we compute two collective variables matching those used in the original D.E. Shaw study: (1) the disulfide torsion angle between Cys14 and Cys38, calculated using the dihedral angle formed by CB14–SG14–SG38–CB38 atoms, and (2) the backbone RMSD of residues 4-54 after mean-centering. Remarkably, our optimized path passes through all intermediate snapshots, indicating that it successfully discovers the same transition mechanism as the computationally intensive explicit-solvent MD simulation.
We matched their simulation conditions by incorporating explicit water molecules in our model to ensure a meaningful comparison with the Anton simulation. We generated water shells extending 14Å from any protein atom for both the initial and final BPTI conformations. To handle water molecules during transitions, we pair proximal water molecules between start and end states, then model their movement using simple parameterized splines, rather than directly incorporating all water coordinates into the neural network. This hybrid approach allows us to include explicit solvent effects while keeping computational costs low. The specific details of our water shell construction and matching algorithm are described in Appendix B.2.
The system was modeled using the AMBER99sb-ILDN force field (Lindorff-Larsen et al., 2010) with the TIP3P water model (Jorgensen et al., 1983; Neria et al., 1996). Each minimum-energy path was computed in approximately 15 minutes on a single A6000 GPU, compared to the D.E. Shaw simulation, which required approximately 1.3 million node-hours on their specialized Anton supercomputer to generate the 1.03 millisecond trajectory capturing the same conformational changes. Appendix B.2.6 provides a detailed computational performance comparison.
5 Limitations
While our feedforward neural architecture can model conformational transitions, it could benefit from incorporating domain-specific structural priors to improve training efficiency and convergence speed. Physical force fields introduce challenges through high-energy spikes that can cause training instability and divergence in some runs. Our gradient clipping approach mitigates this issue but doesn’t universally prevent training failures.
The method has been validated primarily on small to medium-sized protein systems, and scaling to larger multi-domain proteins will require additional optimization strategies. Our approach is also constrained to systems with differentiable force fields, but this limitation diminishes as more implementations become available.
Finally, while computationally efficient, our method focuses on identifying minimum-energy paths rather than sampling the full statistical ensemble of transition pathways, representing a trade-off between efficiency and mechanistic detail.
6 Conclusion
We introduced a novel approach to molecular transition path discovery that leverages physics-informed neural networks and differentiable force fields to enable continuous representation of conformational changes. By reformulating minimum-energy path discovery as a neural optimization problem, we demonstrated the method’s effectiveness on both the small AIB9 peptide and the explicitly solvated BPTI protein, showing its ability to discover physically meaningful transition pathways across different complexity scales.
Our approach achieves remarkable computational efficiency, reducing computational requirements by approximately six orders of magnitude compared to traditional MD simulations for the BPTI system. These advances make previously intractable conformational transition studies accessible on standard computational resources.
While our current implementation shows promising results, several directions for improvement remain. A key limitation is the simple feedforward architecture used to represent the transition paths. Future work should explore neural architectures that better respect the structural organization of protein systems. Ramakrishnan et al. (2025) investigates non-uniform sampling schemes along the path, among other innovative design decisions that might further aid convergence. Integrating pretrained generative models suggested by Raja et al. (2025) could provide better initial transition guesses, potentially further accelerating convergence by starting optimization from more plausible conformational pathways. Extending the method to handle larger multi-domain proteins is also a promising frontier. Future work could also explore methods to estimate free energies and reaction rates from our discovered paths, which would provide valuable kinetic and thermodynamic insights beyond the current geometric focus.
7 Broader Impacts
Our research democratizes access to molecular simulation by enabling transition path studies on standard hardware. This may accelerate drug discovery and protein engineering while contributing to sustainable computing through drastically reduced energy consumption. We see no direct negative societal impacts, though we acknowledge that accelerated protein engineering capabilities could potentially be misused in the future, and we encourage responsible use with appropriate experimental validation.
Appendix A Technical Appendices and Supplementary Material
A.1 Extended derivation of the energy-based loss function
A.1.1 From Onsager-Machlup to a geometric line integral
We begin with a molecular system evolving under overdamped Langevin dynamics, a standard model for high-friction molecular motion:
where represents the molecular configuration, is the potential energy function, and is a Gaussian white noise with .
The probability density of a trajectory on is given by:
where is the Onsager-Machlup functional:
Expanding the integrand:
For paths with fixed endpoints and , the cross-term integrates to a constant difference in potential energy:
We can define a modified action functional that, when minimized, is equivalent to minimizing the original Onsager-Machlup functional for fixed endpoints:
By the Cauchy-Schwarz inequality:
with equality when is parallel or antiparallel to . This is precisely the defining characteristic of the minimum-energy path.
When this equality holds, our action simplifies to:
The term represents the infinitesimal arc length element along the path, allowing us to rewrite the action as a line integral:
where is the path in configuration space. This geometric form is independent of the path’s parameterization and depends solely on the potential energy landscape.
A.1.2 Discretization and connection to energy differences
To compute this geometric action numerically, we introduce a progress coordinate and a mapping such that and . The relationship between the progress coordinate and arc length is:
We discretize the path with points, setting and for . The discretized geometric action becomes:
where .
Using Taylor’s theorem and assuming , we can relate energy differences to gradient magnitudes:
where is a remainder of order and is the angle between the gradient and displacement vectors.
For a path following the minimum-energy trajectory, the gradient aligns with the path direction ( as ), giving:
This yields an approximation of the geometric action:
A.1.3 Establishing the energy-based loss function
We need to bound the sum of absolute energy differences to establish a direct connection to an energy-based loss function. Using the triangle inequality, for any two points and any constant :
Setting and summing over all segments:
We can rewrite the first sum as and the second as . Combining these terms:
Since , , and are constants, minimizing is equivalent to minimizing the upper bound on the geometric action. This provides the theoretical justification for our energy-based loss function.
A.1.4 Neural network implementation
For practical optimization, we parametrize the path using a neural network with parameters . We employ stochastic sampling of the progress variable to construct our loss function:
where are uniformly sampled progress coordinates.
This stochastic approach provides an unbiased estimator of the expected energy along the path:
By minimizing through gradient-based optimization, the neural network learns to represent low-energy pathways connecting the stable states, effectively discovering the minimum-energy path predicted by the Onsager-Machlup framework.
A.2 Connection between Doob’s Lagrangian and Minimum-Energy Path Optimization
We elaborate on the theoretical connections between our energy-based optimization approach and the framework of Doob’s h-transform for sampling transition paths (Du et al., 2024). This comparison is particularly relevant as both methods employ neural networks to represent molecular state transitions—either as a minimum-energy path in our case or the full transition path ensemble in theirs. Notably, the parametrization of the mean trajectory in Du et al. resembles our MEP representation, differing primarily in our use of blending functions and spatial embeddings for the neural network component.
Doob’s Lagrangian formulation considers the following action functional:
This functional quantifies the cost of controlling a stochastic process to achieve desired endpoint conditions, subject to the system’s energy landscape. Intuitively, it measures the amount of "effort" needed to steer the dynamics from starting state A to target state B, with smaller values indicating more probable transition paths.
In this formulation, is the probability density at time given boundary conditions, parameterized as a Gaussian , and is the control vector field satisfying , with representing the reference drift for overdamped dynamics.
This action functional describes a Wasserstein Lagrangian flow as formalized in (Neklyudov et al., 2023) and (Neklyudov et al., 2024). In this framework, one can learn stochastic dynamics from samples, thereby learning a process that defines a time-dependent density evolving from initial to final states.
The key difference with our PINN-MEP approach is that we minimize the action of a single deterministic path rather than a full distribution of paths. Mathematically, this corresponds to working with a single trajectory through configuration space rather than an ensemble of possible trajectories. This simplification is reflected in our loss function, which contains one fewer expectation than the full Doob’s Lagrangian.
A.2.1 Comparison with Doob’s Lagrangian Method
We provide a direct comparison between our PINN-MEP approach and the method of Du et al. in terms of computational efficiency, as evaluated by the number of force field/energy function calls, and the maximum energies observed in the paths, corresponding to the peak of the potential barrier. To provide a rigorous comparison, we conducted benchmarks using experimental conditions identical to those reported by Du et al. for alanine dipeptide: AMBER14 force field, 300K temperature. Table 1 presents these results alongside those reported in Du et al. (2024) and their recent follow-up work lee_scaling_2025. The only exception is that we do not evaluate trajectory probability, as the assumption of optimization over a non-physical path length makes the calculation of this quantity impossible.
| Method | States | # Evaluations | Max Energy | MinMax Energy |
| MCMC variable length | CV | 21.02M | 740.70 695.79 | 52.37 |
| MCMC* | CV | 1.29B | 288.46 128.31 | 60.52 |
| MCMC variable length | relaxed | 187.54M | 412.65 334.70 | 26.97 |
| MCMC | relaxed | 10B | N/A | N/A |
| MCMC variable length | exact | 10B | N/A | N/A |
| MCMC | exact | 10B | N/A | N/A |
| Doob’s Cartesian | exact | 38.40M | 726.40 0.07 | 726.18 |
| Doob’s Cartesian, 2 Mixtures | exact | 51.20M | 709.38 162.37 | 513.72 |
| Doob’s Cartesian, 5 Mixtures | exact | 51.20M | 541.26 278.20 | 247.96 |
| Doob’s Internal | exact | 38.40M | -14.62 0.02 | -14.67 |
| Doob’s Internal, 2 Mixtures | exact | 51.20M | -15.38 0.14 | -15.54 |
| Doob’s Internal, 5 Mixtures | exact | 51.20M | -15.50 0.31 | -15.95 |
| Doob’s Seq2Seq | exact | - | 1505.6 0.45 | 245.05 0.02 |
| w/ Temperature Annealing | exact | - | 1583.18 0.3 | 592.13 0.26 |
| w/ Interatomic Interpolation | exact | - | 1601.57 0.53 | 3.46 0.03 |
| MaxLL | exact | - | 1532.45 | 615 |
| w/ Temperature Annealing | exact | - | 1599.03 | 233 |
| w/ Interatomic Interpolation | exact | - | 1545.32 | 619 |
| PINN-MEP | exact | 15700 | -67.22 3.55 | -70.78 |
PINN-MEP requires approximately 3261× fewer energy function evaluations (15.7K vs. 51.20M) compared to the Doob’s Lagrangian approach while achieving a 4.44 times lower maximum energy barrier. The lower number of required evaluations is likely due to the computation of the loss, which requires sampling their path over both , as we do as well, but also the width of the path. Thus, the tradeoff of sacrificing full ensemble characterisation is expected to come with computational efficiency.
Appendix B Comparison with Other Related Methods
While our primary benchmark comparison focuses on the Doob’s Lagrangian method due to similarity in task and system setup, examining how our approach relates to other methods for transition path discovery/minimum energy path generation is also informative. Here we present results from two related approaches in the literature, though we emphasize that direct numerical comparisons should be treated with caution due to differences in force fields, optimization objectives, and computational frameworks.
| System | Method | Total energy evals | TS energy (eV) |
| Alanine dipeptide | NEB | 8063 | 11.0014 |
| INR | 3046 | 0.4070 |
| Method | CVs | # FF Evals / Path | Runtime / Path |
| MCMC (Two-Way Shooting) | No | 1B | 100 hours |
| Metadynamics | Yes | 1M | 10 hours |
| OM Opt. (Diffusion Model) | No | 10K | 50 min |
Table 2 shows results from Ramakrishnan et al. (2025) comparing Nudged Elastic Band (NEB) with their Implicit Neural Representation (INR) approach for minimum-energy path discovery. Both our PINN-MEP approach and their INR method utilize lightweight, preexisting molecular dynamics force fields, though we employ different such force fields. Their approach requires a more complex loss function that evaluates both energy and forces along with path parameterization and its gradient (path velocity). In contrast, Table 3 presents results from Raja et al. (2025), who take a fundamentally different approach by repurposing pre-trained generative models to perform sampling of transition paths. A key advantage of their method is that the generative model provides significantly better initial guesses for transition paths based on learned molecular distributions. However, their approach requires evaluating both the pretrained model (as a surrogate force field) and its gradient for their loss function. Our PINN-MEP method stands out for its simplicity, using only the energy potential in the loss function without requiring additional gradient terms, while maintaining comparable computational efficiency. These tables corroborate our findings in the Doob’s Lagrangian comparison: methods optimizing for a single minimum-energy path consistently require orders of magnitude fewer force field evaluations than approaches that characterize the full transition path ensemble.
B.0.1 Network Architecture
Model type: Transition MLP with sinusoidal time embedding
Network depth: 10 hidden layers
Network width: 8 width factor (scaled relative to the number of atoms)
Time embedding: Sinusoidal position encoding with maximum period of 100,000
Activation function: ReLU
B.0.2 Training Parameters
Optimizer: Adam
Learning rate: 0.0004
Batch size: 1
Frames per batch: 8
Spring constant (): 0.0
Interpolation method: Linear
B.1 Ablation Studies
This section presents a series of ablation studies examining the impact of various components of our PINN-MEP approach. These experiments help validate our design choices and provide insights into the importance of different architectural and training elements. We conduct these ablation studies on the AIB9 system because Alanine dipeptide, as a benchmark system, is too simple to provide meaningful insights into the performance of our method. For each ablation study, we report the lowest maximum energy of the path, the mean and standard deviation of the energy, and the mean number of force field evaluations to reach the lowest maximum energy. We run each setting in the ablations five times for a maximum of 150000 batches, with a batch size of 1 and 32 frames per batch. For experiments without a spring constant between adjacent sampled frames, the batch size and frames are treated identically. Due to the nature of physical force fields, slight deviations in the path can lead to large changes in energy, leading to similar settings having very different energies. This can limit the ability to compare the performance of different settings. However, we find that the results are still directionally informative.
B.1.1 Embedding Strategy
We compared sinusoidal time embeddings with different maximum periods against simple linear projections of the progress variable. Table 4 shows sinusoidal embeddings with a period of generally outperformed other configurations.
| Embedding Type | Max Period | Mean Energy | Best Energy | Evaluations |
| Sinusoidal | 5832.29 7006.09 | -33.80 | 142400 | |
| Sinusoidal | 4939.20 2836.56 | 1388.48 | 122100 | |
| Sinusoidal | 9508.61 7297.49 | 1273.18 | 106900 | |
| Sinusoidal | 5851.25 4618.35 | 819.25 | 129600 | |
| Linear (None) | – | 9567.01 7578.20 | 303.63 | 127600 |
B.1.2 Spring Regularization
We tested the effect of chain-of-states-like spring constant regularization, as shown in Table 5. Interestingly, we found that a small amount of spring regularization () provided the best results, suggesting that a minimal amount of explicit path continuity enforcement can be beneficial when combined with our neural representation, however, this is not a requirement for the method to work and the additional hyperparameter does not necessarily be worth including.
| Spring Constant | Mean Energy | Best Energy | Evaluations |
| 368.78 416.68 | -278.09 | 130300 | |
| 8862.90 9138.59 | -155.85 | 120600 | |
| 0.0 | 8200.08 15576.30 | 240.85 | 116500 |
| 1156.70 428.50 | 791.59 | 120400 | |
| 7440.13 3339.87 | 4657.28 | 92300 | |
| 45567.30 5015.78 | 40609.10 | 64700 |
B.1.3 Network Architecture
We investigated the effect of network depth and width on transition path quality. Table 6 shows that an 8-layer network with a width factor of 4 achieved the best balance between expressivity and computational efficiency, with wider networks also performing well.
| Depth | Width Factor | Mean Energy | Best Energy | Evaluations |
| 8 | 4 | 2746.08 5381.57 | -168.17 | 112300 |
| 8 | 6 | 6337.10 6284.13 | -278.09 | 79300 |
| 8 | 2 | 10396.20 14039.40 | -278.09 | 113500 |
| 6 | 6 | 6189.19 7667.11 | 334.22 | 112800 |
| 6 | 4 | 9371.15 15277.50 | 240.85 | 111800 |
| 6 | 2 | 9872.01 14229.30 | 152.36 | 122500 |
| 4 | 6 | 13242.00 8432.26 | 6013.58 | 94000 |
| 4 | 4 | 16300.90 5768.37 | 8295.52 | 132100 |
| 4 | 2 | 24916.70 14933.30 | 12697.60 | 78700 |
B.2 BPTI Modeling and Computational Details
B.2.1 System Preparation
As reference conformations, we used the structures from the D.E. Shaw Research BPTI simulation of BPTI (Shaw et al., 2010). The initial and final states for our minimum-energy path optimization were selected as the second snapshot (index 1) and the second-to-last snapshot (index -2) from the provided reference structures, as they were the most separated in the CVs determined in the original study, named the backbone RMSD and the disulfide torsion angle.
B.2.2 Explicit Solvent Shell Construction
For the BPTI system, we developed a sophisticated approach to handle explicit water molecules efficiently:
Initial water box generation: Both start and end configurations were solvated with TIP3P water molecules using a padding distance of 2.0 nm and energy minimized for 15,000 steps.
Water shell extraction: Rather than using the entire water box, we extracted spherical shells of water molecules within a 1.4 nm radius of any protein atom for both the start and end configurations. This significantly reduced the system size while maintaining the crucial first hydration shell.
Radius optimization: We employed a binary search algorithm to determine optimal shell radii that would capture an identical number of water molecules in both start and end configurations, ensuring consistent system sizes throughout the transition.
Water molecule matching: We applied the Hungarian algorithm to optimally pair water molecules between the start and end states based on spatial proximity, minimizing the overall displacement of water molecules during the transition, which ensures that the intial guess for the MEP does not have very large energies stemming from overlapping molecules.
B.2.3 Hybrid Neural Network Architecture
For the BPTI system with explicit solvent, we developed a specialized hybrid transition model to handle the larger system size efficiently:
Protein representation: Neural network with sinusoidal time embedding for the protein atoms (58 residues, 892 atoms)
Water representation: Parameterized spline interpolation between matched water molecules
Control points: 5 control points for the spline interpolation
Water weight: 0.1 (relative weight of water molecule contributions to the loss)
Clamp threshold: (minimum/maximum energy clipping threshold)
B.2.4 Training Parameters
The BPTI MEP was optimized using the following training parameters:
Optimizer: Adam
Learning rate:
Batch size: 1
Frames per batch: 16
Total iterations: 30,000
Random seed: 1
B.2.5 Force Field Parameters
We used the following force field parameters to match the D.E. Shaw simulation conditions:
Protein force field: AMBER99sb-ILDN (amber99sbildn.xml)
Water force field: TIP3P (tip3p.xml)
B.2.6 Computational Performance
For the BPTI system with explicit solvent (approximately 11,000 atoms total):
Training time: Approximately 15 minutes on a single NVIDIA A6000 GPU
Force field evaluations: Approximately 480,000 evaluations (30,000 iterations x 16 frames per batch)
Trajectory generation: 64 frames for the final MEP visualization and analysis
This computational effort is approximately six orders of magnitude less, in node-hours, than the original D.E. Shaw Research simulation on their specialized Anton supercomputer, which required approximately 1.3 million node-hours to generate the 1.03 millisecond trajectory capturing the same conformational changes.

### v4

@@
 Abstract
-Characterizing conformational transitions in physical systems remains a fundamental challenge in the computational sciences.
-Traditional sampling methods like molecular dynamics (MD) or MCMC often struggle with the high-dimensional nature of molecular systems and the high energy barriers of transitions between stable states.
-While these transitions are rare events in simulation timescales, they often represent the most biologically significant processes, for example, the conformational change of an ion channel protein from its closed to open state, which controls cellular ion flow and is crucial for neural signaling.
-Such transitions in real systems may take milliseconds to seconds but could require months or years of continuous simulation to observe even once.
+Characterizing conformational transitions in physical systems remains a fundamental challenge, as traditional sampling methods struggle with the high-dimensional nature of molecular systems and high energy barriers between stable states.
+These rare events often represent the most biologically significant processes, like ion channel opening that controls cellular signaling, yet may require months of continuous simulation to observe.
 We present a method that reformulates transition path generation as a continuous optimization problem solved through physics-informed neural networks (PINNs) inspired by string methods for minimum-energy path (MEP) generation.
-By representing transition paths as implicit neural functions and leveraging automatic differentiation with differentiable molecular dynamics force fields, our method enables the efficient discovery of physically realistic transition pathways without requiring expensive path sampling.
-We demonstrate our method’s effectiveness on two proteins, including an explicitly hydrated bovine pancreatic trypsin inhibitor (BPTI) system with over 8,300 atoms.
-Our approach reproduces the conformational change captured in a landmark millisecond-scale explicit-solvent MD simulation (Shaw et al., 2010), while achieving remarkable computational efficiency, requiring only force field evaluations compared to the approximately 412 billion evaluations in the original study.
-This represents a nearly six orders of magnitude reduction, allowing us to generate the transition pathway in just 15 minutes on a standard GPU rather than weeks on specialized hardware.
+By representing paths as implicit neural functions and leveraging automatic differentiation with differentiable molecular force fields, our method efficiently discovers physically realistic transition pathways without expensive sampling.
+We demonstrate our approach in two proteins, including an explicitly hydrated BPTI system with more than 8,300 atoms.
+Our method identifies a representative minimum-energy path for the conformational change captured in a millisecond-scale MD simulation (Shaw et al., 2010), while requiring only force field evaluations compared to approximately 412 billion in the original study.
+Though our method yields a single pathway rather than an ensemble of transition trajectories, this represents a six orders of magnitude reduction in computational cost, generating a physically meaningful transition in just 15 minutes on a standard GPU rather than weeks on specialized hardware.
 1 Introduction
@@
 The search for efficient methods to identify transition pathways between stable states has a rich history in computational physics.
-Early theoretical work established key principles through the Freidlin-Wentzell (Freidlin & Wentzell, 1998) and Onsager-Machlup (Onsager & Machlup, 1953) functionals, which connect the most probable transition paths to minimizers of specific action functionals in different limiting regimes.
+Early theoretical work established key principles through the Freidlin-Wentzell (Freidlin & Wentzell, 1998) and Onsager-Machlup (Onsager & Machlup, 1953) functionals, which connect the most probable trajectories, including transitions, to minimizers of specific action functionals.
 Building on these foundations, the minimum-energy path (MEP) between stable molecular states represents the most probable transition pathway in the limit of high friction.
@@
 Concurrent advances in scientific machine learning and molecular simulation have created opportunities for novel approaches to this challenge.
-Physics-informed neural networks (PINNs) have revolutionized our ability to solve complex physical problems by embedding differential equations directly into neural architectures (Raissi et al., 2019).
+Physics-informed neural networks (PINNs) have increased our ability to solve complex physical problems by embedding differential equations directly into neural architectures (Raissi et al., 2019).
 Additionally, the development of differentiable molecular force fields (Wang et al., 2023) has enabled seamless integration of physical forcefield models used in MD simulations with modern machine learning frameworks, allowing efficient computation of derivatives needed for path optimization.
-The last decade has seen numerous approaches emerge to tackle the transition path sampling problem.
-Traditional shooting and other sampling methods (Jung et al., 2023; Lazzeri et al., 2023; Bolhuis et al., 2002; Dellago et al., 1998; Laio & Parrinello, 2002; Torrie & Valleau, 1977) rely on expensive molecular dynamics simulations, require initial MD-harvested trajectories to begin with or preexisting knowledge of the transition dynamics, and face inherent tradeoffs between rejection rates and path diversity.
+The last decades have seen numerous approaches emerge to tackle the transition path sampling problem.
+Traditional shooting and other sampling methods (Jung et al., 2023; Lazzeri et al., 2023; Bolhuis et al., 2002; Dellago et al., 1998; Laio & Parrinello, 2002; Torrie & Valleau, 1977) rely on expensive molecular dynamics simulations, require initial MD-harvested trajectories to begin with or preexisting knowledge of the transition dynamics.
 More recent machine learning approaches have explored several directions: variational approaches using Doob’s h-transform (Du et al., 2024; Das & Limmer, 2019; Singh & Limmer, 2023), stochastic control formulations (Yan et al., 2022; Holdijk et al., 2023), and even attempts to sample full transition trajectories directly using diffusion models (Petersen et al., 2023; Han et al., 2024; Jing et al., 2024).
@@
 While each approach offers valuable insights into neural network-based path discovery, our method focuses on a streamlined optimization scheme that efficiently scales to large biomolecular systems with explicit solvent while maintaining fast training times.
-Here, we introduce an approach focused on efficient minimum-energy path discovery as a foundation for understanding molecular transition mechanisms.
-Our method represents transition paths as continuous, differentiable neural functions that map a progress coordinate to molecular configurations.
-This representation, inspired by recent advances in neural implicit representations (Mildenhall et al., 2020; Sitzmann et al., 2020), naturally incorporates physical constraints through a physics-informed neural network (PINN) framework.
-By directly targeting the minimum-energy path through energy minimization, we achieve remarkable computational efficiency, enabling applications to complex biomolecular systems in minutes on standard hardware while extracting the most physically relevant information about transition mechanisms.
+Drawing inspiration from neural implicit representations in computer vision (Mildenhall et al., 2020; Sitzmann et al., 2020), we reformulate the traditional minimum-energy path finding problem using a continuous optimization framework.
+Our approach leverages the universal function approximation capabilities of neural networks to represent conformational transitions as differentiable mappings from a progress coordinate to molecular configurations.
+By embedding physical principles directly into the optimization objective through a physics-informed neural network (PINN) framework, we eliminate the need for explicit reparameterization steps that burden traditional methods.
+This design enables applications to complex biomolecular systems in minutes on standard hardware while extracting physically meaningful transition mechanisms.
 The resulting pathways can serve as valuable starting points for established transition path sampling methods (Dellago et al., 1998; Bolhuis et al., 2002) when exploration of the complete ensemble is desired.
@@
 2.1 Derivation of the Energy-Based Loss Function and Neural Implementation
-To our loss function, we lean on alternative formulations of the Onsager-Machlup action functional as derived by Vanden-Eijnden & Heymann (2008); Olender & Elber (1997) in the context of the string methods.
-We begin with a molecular system evolving under overdamped Langevin dynamics, a well-established model for molecular motion in the high-friction regime characteristic of biomolecular systems:
+To derive our loss function, which when minimized yields a continuous curve representing the minimum-energy path connecting fixed endpoints and , we build upon alternative formulations of the Onsager-Machlup action functional as explored by Vanden-Eijnden & Heymann (2008) and Olender & Elber (1997) in the context of string methods.
+However, we further simplify the approach after reaching the geometric formulation.
+We begin with a molecular system evolving under overdamped Langevin dynamics, a model for molecular motion in the high-friction regime characteristic of biomolecular systems:
 represents the potential energy function, is Gaussian white noise modeling thermal fluctuations, and physical constants are absorbed into the nondimensionalization.
@@
 where are independent standard Gaussian random variables. The conditional probability of transitioning from to follows a Gaussian distribution:
-To obtain the probability of an entire continuous trajectory, we take the product of these conditional probabilities over all time steps and then take the limit as . This yields the path probability:
-The exponent term can be recognized as the Onsager-Machlup action functional, which is defined as:
+Taking the product of these conditional probabilities over all time steps and then taking the limit as yields the path probability:
+The exponent term can be recognized as the Onsager-Machlup action functional Onsager & Machlup (1953):
 Thus, maximizing the probability of a transition path (subject to fixed endpoints and ) is equivalent to minimizing this action functional.
-The action penalizes paths that deviate from following the deterministic drift .
 To further simplify, we expand the integrand of the action functional:
-Applying the Cauchy-Schwarz inequality to the kinetic and potential terms, we can establish an important relationship:
-with equality holding if and only if is parallel (or antiparallel) to .
-When the path is reparameterized so that velocity consistently aligns with the gradient, the action is minimized, yielding the most likely transition pathway.
-For such pathways where is everywhere proportional to , we can transform the time integral into a purely geometric one.
-However, to maintain this parallel alignment, the path length itself must be subject to optimization, as a fixed-length path would generally not satisfy the parallelity condition.
-By defining the arc-length element along the spatial curve traced out by , we obtain:
-This geometric reformulation connects directly to the concept of a minimum-energy path (MEP), which follows the potential energy landscape’s gradient while avoiding high-energy regions.
-The MEP represents the most probable transition mechanism in the overdamped limit, making it an ideal target for our optimization.
-We represent the transition path as a continuous neural mapping for such that and , where represents the network parameters.
-To transform the geometric line integral into a practical loss function, we first discretize the path and apply Taylor expansion to relate gradient magnitudes to energy differences:
-where and . Using the triangle inequality, we can establish that minimizing provides an upper bound on this action, which yields our neural network loss function:
+For paths with fixed endpoints, the cross-term integrates to a constant difference in potential energy:
+Since and are fixed in our formulation, this term becomes a constant that doesn’t affect the optimization.
+We can therefore focus on minimizing the remaining terms.
+Applying the Cauchy-Schwarz inequality:
+This upper bound becomes an equality when is parallel or antiparallel to , with the direction determined by whether the path is ascending or descending the energy landscape.
+Crucially, in our formulation, we allow the path to optimize its shape and effective length.
+This freedom enables the path to adjust to follow the energy gradient (or its opposite), making the above upper bound tight at the optimum.
+When we abandon fixed-time parameterization and allow the path to stretch as needed to align with energy gradients, we can transform the time integral into a purely geometric one by defining the arc-length element along the spatial curve :
+[…]
+where and .
+For a path segment between adjacent points, applying Taylor’s theorem, we can express the energy difference as:
+Where is the angle between the gradient and displacement vectors. For minimum-energy paths, the displacement aligns with or against the gradient direction () giving:
+This gives us a surrogate objective with the same minimum:
+Using the triangle inequality on the absolute energy differences, we can establish that minimizing the simple sum of energies along the path provides an upper bound on this action:
+Since the latter terms are constants during optimization, this leads to our final loss function.
+To implement this in practice, we sample B points along a parametrized path and directly minimize the sum of energies at these points, yielding:
+where are uniformly or randomly sampled points in .
 A complete derivation with all mathematical details is provided in Appendix A.1.
@@
 2.1.1 Neural Representation of Minimum-Energy Paths
-Drawing inspiration from the string method formalism, we propose representing the transition path using an implicit neural parameterization, similar to the mean parameterization of Du et al.
+We propose representing the path using a neural parameterization, similar to the mean parameterization of Du et al.
 (Du et al., 2024).
@@
 Specifically, we construct a continuous function mapping a progress coordinate to molecular configurations , where is the number of atoms:
-Here is a neural network with parameters , and are basis functions that satisfy the endpoint constraints , , , and , ensuring that and regardless of the network’s output.
+Here is a neural network, and are basis functions that satisfy the endpoint constraints , , , and , ensuring that and regardless of the network’s output.
 While simple linear functions and can be used, other basis functions with appropriate boundary conditions are also effective.
@@
 2.1.2 Path Optimization
-Rather than alternating between evolution and reparameterization steps as in the classical string method, we minimize directly through stochastic gradient descent. For each batch of paths:
-The implicit neural representation naturally handles reparameterization constraints through its continuous parameterization.
-Unlike traditional chain-of-states methods, explicit spring-like continuity terms are unnecessary—the inherent continuity of , combined with smooth blending functions, ensures path continuity without additional constraints.
+We minimize directly through stochastic gradient descent. For each batch of points along the path:
+Unlike traditional chain-of-states methods, explicit spring-like continuity terms are unnecessary: the inherent continuity of , combined with smooth blending functions, ensures path continuity without additional constraints.
 2.2 Established Approaches to Minimum-Energy Path Finding
@@
 2.2.1 String Method
-The string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Ramakrishnan et al., 2025) determines MEPs by evolving a curve in configuration space according to physical forces while maintaining proper parameterization.
+The string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Ramakrishnan et al., 2025) determines the MEPs by evolving a curve in the configuration space according to the physical forces while maintaining proper parameterization.
 Given the same endpoints and used in our formulation, the method seeks a curve (string) parameterized by that connects these states.
@@
 This two-step procedure converges to a discretized MEP satisfying each image’s local orthogonality condition .
-Unlike the neural approach presented in this paper, which handles parameterization implicitly, this method requires explicit reparameterization steps that can introduce computational inefficiencies.
+This method requires explicit reparameterization steps that can introduce computational inefficiencies.
 2.2.2 Chain-of-States Methods
@@
 where and is a spring constant.
-The spring term enforces continuity of the path while allowing flexibility in the discretization.
+The spring term enforces the continuity of the path while allowing flexibility in the discretization.
 However, this formulation introduces a well-known limitation: the spring forces tend to cause "corner cutting," where the path takes shortcuts through high-energy regions rather than following the low-energy valleys representing physically meaningful transitions.
-The neural representation described earlier avoids this limitation through its natural continuity properties.
-At convergence, both formulations yield a discretized MEP satisfying at each image, the same condition that PINN-MEP targets through energy minimization, but with the advantage of continuous representation and gradient-based optimization.
+Although variants of chain-of-states methods exist that avoid this limitation, they come at an increased computational cost.
+The neural representation avoids this limitation through its natural continuity properties.
 3 Experiments
@@
 These systems represent different scales of complexity and serve to validate our method’s ability to discover physically meaningful transition paths.
-In Appendix A.2, we compare and benchmark our method to the related approach outlined in (Du et al., 2024), and in Appendix B.1, we present comprehensive ablation studies analyzing the impact of different embedding strategies, interpolation methods, network architectures, and regularization approaches.
+In Appendix A.2, we compare and benchmark our method with the related approach outlined in (Du et al., 2024), and in Appendix B.1, we present ablation studies analyzing the impact of different embedding strategies, interpolation methods, network architectures, and regularization approaches.
 3.1 AIB9 Transition Path Discovery
@@
 Bovine Pancreatic Trypsin Inhibitor (BPTI) is a clinically significant protein used as an anticoagulant in medical procedures.
-Beyond its therapeutic applications, BPTI has become a fundamental model system for studying protein dynamics due to its moderate size (58 residues) and complex conformational changes involving disulfide bond rearrangements.
-To investigate BPTI’s conformational transitions, we analyze five key states from a landmark molecular dynamics (MD) simulation performed by D.E.
+BPTI is 58 residues long and has complex conformational changes involving disulfide bond rearrangements paired with backbone changes.
+To investigate BPTI’s conformational transitions, we use as reference structures five key states from a landmark molecular dynamics (MD) simulation performed by D.E.
 Shaw Research on their Anton supercomputer (Shaw et al., 2010).
-These publicly available snapshots provide rare insight into BPTI’s transition mechanisms.
+These publicly available snapshots provide rare benchmark structures for validating our approach to BPTI’s transition mechanisms.
 Full details of our BPTI system preparation, explicit solvent shell construction, and hybrid neural network architecture for handling water molecules are provided in Appendix B.2.
-We generate a transition MEP connecting the initial and final BPTI states from the D.E.
+We generate multiple transition MEPs connecting two BPTI states from the D.E.
 Shaw simulation.
@@
 To handle water molecules during transitions, we pair proximal water molecules between start and end states, then model their movement using simple parameterized splines, rather than directly incorporating all water coordinates into the neural network.
-This hybrid approach allows us to include explicit solvent effects while keeping computational costs low.
+This hybrid neural-spline approach allows us to include explicit solvent effects while keeping computational costs low.
 The specific details of our water shell construction and matching algorithm are described in Appendix B.2.
@@
 Shaw simulation, which required approximately 1.3 million node-hours on their specialized Anton supercomputer to generate the 1.03 millisecond trajectory capturing the same conformational changes.
+While our approach does not provide the same statistical ensemble information as equilibrium MD, it efficiently captures the key transition pathway.
+It’s worth noting that even the reference millisecond-scale simulation may not have fully achieved equilibrium sampling of the conformational space.
 Appendix B.2.7 provides a detailed computational performance comparison.
@@
 While our feedforward neural architecture can model conformational transitions, it could benefit from incorporating domain-specific structural priors to improve training efficiency and convergence speed.
-Physical force fields introduce challenges through high-energy spikes that can cause training instability and divergence in some runs.
+Furthermore, physical force fields introduce challenges through high-energy spikes that can cause training instability and divergence in some runs.
 Our gradient clipping approach mitigates this issue but doesn’t universally prevent training failures.
-The method has been validated primarily on small to medium-sized protein systems, and scaling to larger multi-domain proteins will require additional optimization strategies.
+The method has been validated primarily on small to medium-sized protein systems, and scaling to larger multi-domain proteins could require additional optimization strategies.
 Our approach is also constrained to systems with differentiable force fields, but this limitation diminishes as more implementations become available.
-Finally, while computationally efficient, our method focuses on identifying minimum-energy paths rather than sampling the full statistical ensemble of transition pathways, representing a trade-off between efficiency and mechanistic detail.
+Finally, while computationally efficient, our method focuses on identifying minimum-energy paths rather than sampling the full statistical ensemble of transition pathways, representing a trade-off between efficiency and exhaustively sampling a system.
 6 Conclusion
-We introduced a novel approach to molecular transition path discovery that leverages physics-informed neural networks and differentiable force fields to enable continuous representation of conformational changes.
+We introduced a novel approach to molecular transition path discovery that leverages physics-informed neural networks and differentiable force fields to enable the continuous representation of conformational changes.
 By reformulating minimum-energy path discovery as a neural optimization problem, we demonstrated the method’s effectiveness on both the small AIB9 peptide and the explicitly solvated BPTI protein, showing its ability to discover physically meaningful transition pathways across different complexity scales.
 Our approach achieves remarkable computational efficiency, reducing computational requirements by approximately six orders of magnitude compared to traditional MD simulations for the BPTI system.
-These advances make previously intractable conformational transition studies accessible on standard computational resources.
+While our method yields a single pathway rather than the complete ensemble of transition trajectories, this represents a valuable trade-off that makes previously intractable conformational transition studies accessible on standard computational resources.
 While our current implementation shows promising results, several directions for improvement remain.
-A key limitation is the simple feedforward architecture used to represent the transition paths.
 Future work should explore neural architectures that better respect the structural organization of protein systems.
@@
 Integrating pretrained generative models suggested by Raja et al.
-(2025) could provide better initial transition guesses, potentially further accelerating convergence by starting optimization from more plausible conformational pathways.
+(2025) could provide better initial transition guesses, potentially accelerating convergence by starting optimization from more plausible conformational pathways.
 Extending the method to handle larger multi-domain proteins is also a promising frontier.
-Future work could also explore methods to estimate free energies and reaction rates from our discovered paths, which would provide valuable kinetic and thermodynamic insights beyond the current geometric focus.
+Future work could also explore methods to estimate free energies and reaction rates from our discovered paths.
 7 Broader Impacts
-Our research democratizes access to molecular simulation by enabling transition path studies on standard hardware.
+Our research democratizes access to the study of larger molecular systems by enabling transition path studies on standard hardware.
 This may accelerate drug discovery and protein engineering while contributing to sustainable computing through drastically reduced energy consumption.
-We see no direct negative societal impacts, though we acknowledge that accelerated protein engineering capabilities could potentially be misused in the future, and we encourage responsible use with appropriate experimental validation.
+We see no direct negative societal impacts, though we acknowledge that accelerated protein engineering capabilities could be misused in the future, and we encourage responsible use with appropriate experimental validation.
 Appendix A Technical Appendices and Supplementary Material
 A.1 Extended derivation of the energy-based loss function
+In this appendix, we provide a comprehensive derivation of our energy-based loss function for discovering minimum-energy paths (MEPs) between molecular configurations.
+The derivation proceeds in four main steps: (1) establishing the path probability density from overdamped Langevin dynamics, (2) reformulating this probability in terms of the Onsager-Machlup action functional, (3) simplifying this functional to a geometric line integral, and (4) developing a discretized approximation suitable for neural network optimization.
+Our goal is to find the most probable transition path between two stable molecular configurations and .
+This path corresponds to the minimum-energy pathway through the potential energy landscape and represents the most likely mechanism of conformational change.
+Assumptions and Approximations
+The derivation relies on several key assumptions:
+The molecular system evolves according to overdamped Langevin dynamics, appropriate for high-friction biomolecular environments
+The path endpoints and are fixed and correspond to stable states
+The potential energy function is differentiable everywhere along the path
+We can represent the path as a continuous, parameterized curve in configuration space with sufficient flexibility
+The path’s parameterization allows for effective length changes, permitting the path to adjust its stretching as needed to follow energy gradients optimally
+We also make several important approximations during the derivation:
+We use a first-order Taylor expansion to relate energy differences to gradient magnitudes
+We introduce stochastic sampling to estimate the continuous energy integral
+For a minimum-energy path, the displacement tends to align either parallel or anti-parallel with the gradient direction (i.e., for path segments)
+We use the triangle inequality to establish an upper bound on the sum of absolute energy differences
+While the last two approximations do not change the location of the global minimum, they do affect the optimization landscape before that minimum is reached.
 A.1.1 From Onsager-Machlup to a geometric line integral
@@
 where represents the molecular configuration, is the potential energy function, and is a Gaussian white noise with .
-The probability density of a trajectory on is given by:
-where is the Onsager-Machlup functional:
+To derive the path probability, we discretize the time interval into steps of duration , approximating the dynamics as:
+where are independent standard Gaussian random variables with zero mean and unit variance.
+The conditional probability of transitioning from to follows a Gaussian distribution:
+where is the dimensionality of the system. The probability of the entire discretized path is given by the product of these conditional probabilities:
+Taking the limit as and with fixed, this sum approaches a path integral:
+Thus, the path probability can be written as:
+The exponent term is recognized as the Onsager-Machlup action functional:
+This action functional penalizes paths that deviate from following the force .
+Maximizing the probability of a transition path (subject to fixed endpoints and ) is equivalent to minimizing this action functional.
 Expanding the integrand:
@@
 By the Cauchy-Schwarz inequality:
-with equality when is parallel or antiparallel to . This is precisely the defining characteristic of the minimum-energy path.
+with equality when is parallel or anti-parallel to , with the direction determined by whether the path is ascending or descending the energy landscape.
+This is precisely the defining characteristic of the minimum-energy path (MEP) - a path that follows the potential energy landscape’s gradient while avoiding high-energy regions.
 When this equality holds, our action simplifies to:
@@
 A.1.2 Discretization and connection to energy differences
-To compute this geometric action numerically, we introduce a progress coordinate and a mapping such that and . The relationship between the progress coordinate and arc length is:
+To compute this geometric action numerically, we introduce a progress coordinate and a mapping such that and , where represents the network parameters.
+The relationship between the progress coordinate and arc length is:
 We discretize the path with points, setting and for . The discretized geometric action becomes:
 where .
-Using Taylor’s theorem and assuming , we can relate energy differences to gradient magnitudes:
-where is a remainder of order and is the angle between the gradient and displacement vectors.
-For a path following the minimum-energy trajectory, the gradient aligns with the path direction ( as ), giving:
+Using Taylor’s theorem, we can relate energy differences to gradient magnitudes:
+where is the remainder of the order and is the angle between the gradient and displacement vectors.
+For a path following the minimum-energy trajectory, the gradient aligns with or against the path direction ( as , depending on whether the path is ascending or descending the energy landscape), giving at the minimum of the optimization:
 This yields an approximation of the geometric action:
+This forms a surrogate objective with the same minimum as our original action.
 A.1.3 Establishing the energy-based loss function
 We need to bound the sum of absolute energy differences to establish a direct connection to an energy-based loss function. Using the triangle inequality, for any two points and any constant :
-Setting and summing over all segments:
+Choosing to be the minimum energy along the entire path and summing over all segments:
 We can rewrite the first sum as and the second as . Combining these terms:
-Since , , and are constants, minimizing is equivalent to minimizing the upper bound on the geometric action. This provides the theoretical justification for our energy-based loss function.
+Since , , and are constants during optimization, minimizing is equivalent to minimizing the upper bound on the geometric action.
+This provides the theoretical justification for our energy-based loss function.
 A.1.4 Neural network implementation
-For practical optimization, we parametrize the path using a neural network with parameters . We employ stochastic sampling of the progress variable to construct our loss function:
-where are uniformly sampled progress coordinates.
+For practical optimization, we parametrize the path using a neural network with parameters .
+To implement our approach, we sample points along the path and minimize the energy at these sampled locations.
+This represents a transition from the theoretical derivation using discrete segments to a practical implementation using sampled points for more efficient optimization.
+We employ stochastic sampling of the progress variable to construct our loss function:
+where are uniformly sampled progress coordinates and is the batch size (the number of points sampled along the path during each optimization step).
 This stochastic approach provides an unbiased estimator of the expected energy along the path:
 By minimizing through gradient-based optimization, the neural network learns to represent low-energy pathways connecting the stable states, effectively discovering the minimum-energy path predicted by the Onsager-Machlup framework.
+The resulting path follows the potential energy landscape’s gradient while avoiding high-energy regions, representing the most probable transition mechanism in the overdamped limit.
 A.2 Connection between Doob’s Lagrangian and Minimum-Energy Path Optimization
@@
 This functional quantifies the cost of controlling a stochastic process to achieve desired endpoint conditions, subject to the system’s energy landscape.
-Intuitively, it measures the amount of "effort" needed to steer the dynamics from starting state A to target state B, with smaller values indicating more probable transition paths.
+Intuitively, it measures the amount of "effort"/action needed to steer the dynamics from starting state A to target state B, with smaller values indicating more probable transition paths.
 In this formulation, is the probability density at time given boundary conditions, parameterized as a Gaussian , and is the control vector field satisfying , with representing the reference drift for overdamped dynamics.
@@
 We provide a direct comparison between our PINN-MEP approach and the method of Du et al.
-in terms of computational efficiency, as evaluated by the number of force field/energy function calls, and the maximum energies observed in the paths, corresponding to the peak of the potential barrier.
+in terms of computational efficiency, as evaluated by the number of force-field/energy function calls, and the maximum energies observed in the paths, corresponding to the peak of the potential barrier.
 To provide a rigorous comparison, we conducted benchmarks using experimental conditions identical to those reported by Du et al.
@@
 Table 1 presents these results alongside those reported in Du et al.
-(2024) and their recent follow-up work lee_scaling_2025.
+(2024) and their recent follow-up work Lee et al.
+(2025).
 The only exception is that we do not evaluate trajectory probability, as the assumption of optimization over a non-physical path length makes the calculation of this quantity impossible.
@@
 These tables corroborate our findings in the Doob’s Lagrangian comparison: methods optimizing for a single minimum-energy path consistently require orders of magnitude fewer force field evaluations than approaches that characterize the full transition path ensemble.
-B.0.1 Network Architecture
-Model type: Transition MLP with sinusoidal time embedding
-Network depth: 10 hidden layers
-Network width: 8 width factor (scaled relative to the number of atoms)
-Time embedding: Sinusoidal position encoding with maximum period of 100,000
-Activation function: ReLU
-B.0.2 Training Parameters
-Optimizer: Adam
-Learning rate: 0.0004
-Batch size: 1
-Frames per batch: 8
-Spring constant (): 0.0
-Interpolation method: Linear
 B.1 Ablation Studies
@@
 | 4 | 2 | 24916.70 14933.30 | 12697.60 | 78700 |
+B.1.4 AIB9 Network Architecture
+Model type: Transition MLP with sinusoidal time embedding
+Network depth: 6 hidden layers
+Network width: 4 width factor (scaled relative to the number of atoms)
+Time embedding: Sinusoidal position encoding with maximum period of 1,000,000
+Activation function: ReLU
+B.1.5 Training Parameters
+Optimizer: Adam
+Learning rate: 0.00001
+Batch size: 4
+Frames per batch: 32
+Spring constant (): 0.0
+Interpolation method: Linear
 B.2 BPTI Modeling and Computational Details
@@
 For the BPTI system with explicit solvent, we developed a specialized hybrid transition model to handle the larger system size efficiently:
-Protein representation: Neural network with sinusoidal time embedding for the protein atoms (58 residues, 892 atoms)
+Protein representation: Neural network with 6 layers, a width factor of 4, and sinusoidal time embedding for the protein atoms (58 residues, 892 atoms)
 Water representation: Parameterized spline interpolation between matched water molecules
-Control points: 5 control points for the spline interpolation
+Control points: 7 control points for the spline interpolation
 Water weight: 0.1 (relative weight of water molecule contributions to the loss)
@@
 Total iterations: 30,000
-Random seed: 1
-B.2.5 Force Field Parameters
+B.2.5 Handling High-Energy Spikes During Training
+Physical force fields present unique optimization challenges due to occasional extreme energy spikes when atoms approach one another too closely.
+To address this, we implemented adaptive gradient clipping in our training procedure:
+where is a clipping threshold (set to in our experiments).
+We specifically use adaptive gradient clipping rather than a fixed threshold because our hybrid architecture contains two fundamentally different parameter types: neural network weights for protein coordinates and spline control points for water molecules.
+The adaptive nature of the clipping automatically adjusts to the appropriate scale for each parameter type.
+Additionally, we apply different scaling factors to these parameter groups:
+with for water-related parameters.
+This differential treatment further balances the influence of each component during optimization.
+While this approach does not completely eliminate training instabilities, we find it sufficient for most protein systems when combined with careful initialization and learning rate scheduling.
+B.2.6 Force Field Parameters
 We used the following force field parameters to match the D.E. Shaw simulation conditions:
@@
 Water force field: TIP3P (tip3p.xml)
-B.2.6 Computational Performance
-For the BPTI system with explicit solvent (approximately 11,000 atoms total):
-Training time: Approximately 15 minutes on a single NVIDIA A6000 GPU
+B.2.7 Computational Performance
+For the BPTI system with explicit solvent (approximately 8,300 atoms total):
+Training time: Approximately 7 to 15 minutes on a single NVIDIA A6000 GPU
 Force field evaluations: Approximately 480,000 evaluations (30,000 iterations x 16 frames per batch)

### v5

@@
 By representing paths as implicit neural functions and leveraging automatic differentiation with differentiable molecular force fields, our method efficiently discovers physically realistic transition pathways without expensive sampling.
-We demonstrate our approach in two proteins, including an explicitly hydrated BPTI system with more than 8,300 atoms.
+We demonstrate our approach on two proteins, including an explicitly hydrated BPTI system with more than 8,300 atoms.
 Our method identifies a representative minimum-energy path for the conformational change captured in a millisecond-scale MD simulation (Shaw et al., 2010), while requiring only force field evaluations compared to approximately 412 billion in the original study.
-Though our method yields a single pathway rather than an ensemble of transition trajectories, this represents a six orders of magnitude reduction in computational cost, generating a physically meaningful transition in just 15 minutes on a standard GPU rather than weeks on specialized hardware.
+Though our method yields a single pathway rather than an ensemble of transition trajectories, this represents a six orders of magnitude reduction in computational cost, generating a physically meaningful transition in less than 15 minutes on a standard GPU rather than weeks on specialized hardware.
 1 Introduction
-Understanding molecular conformational changes is fundamental to understanding biological processes, from enzyme catalysis to protein folding.
+Characterizing molecular conformational changes, or transitions, is essential for understanding biological processes, from enzyme catalysis to protein folding.
 Although molecular dynamics (MD) simulations provide unprecedented atomic-level detail of these processes (Dror et al., 2012), capturing rare transition events remains computationally challenging due to the inherent timescale gap between molecular motion and conformational changes.
 This difficulty arises from transition rates between stable states being exponentially suppressed by the height of the energy barriers that separate them, as described by transition state theory (Eyring, 1935).
-As a result, direct simulation methods spend the vast majority of computational effort sampling stable states, making the observation of transitions between them prohibitively expensive.
+As a result, direct simulation methods spend the vast majority of computational effort sampling already known stable states, making the observation of transitions between them prohibitively expensive.
 The search for efficient methods to identify transition pathways between stable states has a rich history in computational physics.
@@
 Moreover, the MEP can serve as a scaffold for subsequent sampling of the full transition path ensemble, making it a powerful tool for understanding molecular dynamics.
-Methods such as the string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Petersen et al., 2024; Dellago et al., 1998) approximate these paths through a discrete series of molecular configurations, connected by artificial spring forces to ensure continuity or as splines.
-Although grounded in solid theoretical principles, these approaches face fundamental challenges stemming from their optimization in Cartesian molecular coordinates.
+Methods such as the string method family or chain-of-states approaches (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Petersen et al., 2024; Dellago et al., 1998) approximate these paths as splines or as a discrete series of molecular configurations connected by artificial spring forces to ensure continuity.
+Despite their strong theoretical foundation, these approaches face fundamental challenges stemming from their optimization in Cartesian molecular coordinates.
 The discrete chain of conformations requires periodic reparameterization to maintain even spacing, leading to a complex alternation between energy minimization and geometric constraints.
 Additional limitations include the difficulty of generating good initial transition guesses and limited parallelization, which restricts the scale of practical applications.
-While this formulation has proven useful for many small-scale systems, it has not yet fully benefited from modern optimization approaches that could address these challenges.
-Concurrent advances in scientific machine learning and molecular simulation have created opportunities for novel approaches to this challenge.
+While this formulation has proven useful for small-scale systems, it has not yet benefited from modern optimization approaches that could address these challenges.
+Recent advances in machine learning methodology have created opportunities for novel approaches to this challenge.
 Physics-informed neural networks (PINNs) have increased our ability to solve complex physical problems by embedding differential equations directly into neural architectures (Raissi et al., 2019).
-Additionally, the development of differentiable molecular force fields (Wang et al., 2023) has enabled seamless integration of physical forcefield models used in MD simulations with modern machine learning frameworks, allowing efficient computation of derivatives needed for path optimization.
-The last decades have seen numerous approaches emerge to tackle the transition path sampling problem.
-Traditional shooting and other sampling methods (Jung et al., 2023; Lazzeri et al., 2023; Bolhuis et al., 2002; Dellago et al., 1998; Laio & Parrinello, 2002; Torrie & Valleau, 1977) rely on expensive molecular dynamics simulations, require initial MD-harvested trajectories to begin with or preexisting knowledge of the transition dynamics.
-More recent machine learning approaches have explored several directions: variational approaches using Doob’s h-transform (Du et al., 2024; Das & Limmer, 2019; Singh & Limmer, 2023), stochastic control formulations (Yan et al., 2022; Holdijk et al., 2023), and even attempts to sample full transition trajectories directly using diffusion models (Petersen et al., 2023; Han et al., 2024; Jing et al., 2024).
+Additionally, the development of differentiable molecular force fields (Wang et al., 2023) has enabled seamless integration of physical force field models used in MD simulations with modern machine learning frameworks, allowing efficient computation of derivatives beyond forces, which are needed for neural path optimization.
+Numerous approaches have emerged over the past decades to tackle the problem of sampling transitions.
+Traditional simulation-based sampling methods (Jung et al., 2023; Lazzeri et al., 2023; Bolhuis et al., 2002; Dellago et al., 1998; Laio & Parrinello, 2002; Torrie & Valleau, 1977) rely on expensive molecular dynamics simulations, often requiring initial MD-harvested transition trajectories or prior knowledge of the transition dynamics to begin with.
+More recent machine learning approaches have explored several directions: variational approaches using Doob’s h-transform (Du et al., 2024; Lee et al., 2025; Das & Limmer, 2019; Singh & Limmer, 2023), stochastic control formulations (Yan et al., 2022; Holdijk et al., 2023), and even attempts to sample transition trajectories directly using diffusion models (Petersen et al., 2023; Han et al., 2024; Jing et al., 2024).
 Notably, Du et al.
-(2024) use a similar neural representation of transition paths but with a different objective, targeting the complete transition path ensemble via Doob’s h-transform.
-This approach provides rich statistical information about transition mechanisms.
+(2024) uses a similar neural representation of transition paths but with a different objective, targeting the complete transition path ensemble by additionally parameterizing path width and optimizing based on a framework rooted in Doob’s h-transform (Doob, 1957) and Wasserstein Lagrangian flows (Neklyudov et al., 2024).
 Most closely related to our method, Ramakrishnan et al.
-(2025) employ a continuous neural parameterization of minimum-energy paths inspired by Nudged Elastic Band (NEB), incorporating specialized components for tangential forces, path velocity, and transition state energetics.
+(2025) employ a continuous neural parameterization of MEPs inspired by Nudged Elastic Band (NEB) methods, a class of methods related to string methods, incorporating specialized components for tangential forces, path velocity, and transition state energetics to optimize the neural path.
 Another promising concurrent approach by Raja et al.
-(2025) creatively leverages pretrained diffusion models as a sampler of initial transition conformations and as a forcefield within a chain-of-states framework to generate plausible transition distributions.
-While each approach offers valuable insights into neural network-based path discovery, our method focuses on a streamlined optimization scheme that efficiently scales to large biomolecular systems with explicit solvent while maintaining fast training times.
-Drawing inspiration from neural implicit representations in computer vision (Mildenhall et al., 2020; Sitzmann et al., 2020), we reformulate the traditional minimum-energy path finding problem using a continuous optimization framework.
-Our approach leverages the universal function approximation capabilities of neural networks to represent conformational transitions as differentiable mappings from a progress coordinate to molecular configurations.
-By embedding physical principles directly into the optimization objective through a physics-informed neural network (PINN) framework, we eliminate the need for explicit reparameterization steps that burden traditional methods.
-This design enables applications to complex biomolecular systems in minutes on standard hardware while extracting physically meaningful transition mechanisms.
-The resulting pathways can serve as valuable starting points for established transition path sampling methods (Dellago et al., 1998; Bolhuis et al., 2002) when exploration of the complete ensemble is desired.
+(2025) creatively leverages pretrained diffusion models as a sampler of initial transition conformations and as a force field to generate plausible transition paths using the discretized Onsager-Machlup action directly as a loss function.
+While each approach offers valuable insights into neural network-based path discovery, our method focuses on a streamlined optimization scheme that efficiently scales to much larger biomolecular systems with explicit solvent, essential for accurately modeling such systems, while maintaining fast training times.
+Drawing inspiration from neural implicit representations in computer vision (Mildenhall et al., 2020; Sitzmann et al., 2020), we reformulate the traditional minimum-energy-path finding problem as a continuous optimization framework.
+Our approach transforms the Onsager-Machlup functional, which naturally models the probability of paths in stochastic systems, into a remarkably simple and intuitive loss function that requires only energy evaluations along the neural network parametrized path.
+In contrast to existing methods that use complex composite losses involving energies, forces, and path gradients, our streamlined formulation eliminates the need for these additional computations.
+This simplification, while faster, also sidesteps known issues in PINN optimization if the loss contains differential operators (Rathore et al., 2024).
+This PINN framework efficiently captures physically meaningful transition mechanisms for complex biomolecular systems, completing optimization in minutes on standard hardware.
 2 Theory and Methods
 2.1 Derivation of the Energy-Based Loss Function and Neural Implementation
-To derive our loss function, which when minimized yields a continuous curve representing the minimum-energy path connecting fixed endpoints and , we build upon alternative formulations of the Onsager-Machlup action functional as explored by Vanden-Eijnden & Heymann (2008) and Olender & Elber (1997) in the context of string methods.
-However, we further simplify the approach after reaching the geometric formulation.
+To derive our loss function, which when minimized yields a continuous curve representing the MEP connecting fixed endpoints and , we build upon alternative formulations of the Onsager-Machlup action functional as explored by Vanden-Eijnden & Heymann (2008) and Olender & Elber (1997) in the context of string methods.
+However, we further simplify the approach after reaching their geometric formulation.
+A complete derivation with all mathematical details is provided in Appendix A.1.
 We begin with a molecular system evolving under overdamped Langevin dynamics, a model for molecular motion in the high-friction regime characteristic of biomolecular systems:
@@
 Taking the product of these conditional probabilities over all time steps and then taking the limit as yields the path probability:
-The exponent term can be recognized as the Onsager-Machlup action functional Onsager & Machlup (1953):
+The exponent corresponds to the Onsager-Machlup action functional Onsager & Machlup (1953):
 Thus, maximizing the probability of a transition path (subject to fixed endpoints and ) is equivalent to minimizing this action functional.
+Notably, there are different derivations of the action that yield an additional higher-order derivative term related to trajectory entropy, whose necessity is conditional on the task (Adib, 2008).
+Since we are primarily interested in MEPs, we neglect this term.
+However, the resultant MEPs still empirically align well with the free energy landscape.
 To further simplify, we expand the integrand of the action functional:
@@
 We can therefore focus on minimizing the remaining terms.
-Applying the Cauchy-Schwarz inequality:
+Applying the Cauchy-Schwarz inequality to the remaining terms:
 This upper bound becomes an equality when is parallel or antiparallel to , with the direction determined by whether the path is ascending or descending the energy landscape.
-Crucially, in our formulation, we allow the path to optimize its shape and effective length.
+Crucially, we allow the path to optimize its shape and length in our formulation.
 This freedom enables the path to adjust to follow the energy gradient (or its opposite), making the above upper bound tight at the optimum.
 When we abandon fixed-time parameterization and allow the path to stretch as needed to align with energy gradients, we can transform the time integral into a purely geometric one by defining the arc-length element along the spatial curve :
-[…]
+While the referenced works (Vanden-Eijnden & Heymann, 2008; Olender & Elber, 1997) derive and use this geometric action, we further simplify it.
+This geometric integral represents a natural endpoint for theoretical development for traditional chain-of-states and string methods.
+In these classical approaches, forces directly update molecular conformations to minimize , making minimization along a discretized string using forces conceptually straightforward.
+However, we want a loss function that acts directly on the neural network parameters in our neural network formulation.
+While we could parameterize the path using a neural network and then compute forces along the path before backpropagating to our parameters, this would introduce additional gradient computations, which have been noted to lead to optimization challenges in PINNs (Rathore et al., 2024).
+We avoid this issue by simplifying to an energy-based loss.
+We represent the transition path as a continuous neural mapping from a progress coordinate to the system’s Cartesian coordinates, such that and . We then discretize the path into segments:
 where and .
-For a path segment between adjacent points, applying Taylor’s theorem, we can express the energy difference as:
-Where is the angle between the gradient and displacement vectors. For minimum-energy paths, the displacement aligns with or against the gradient direction () giving:
+For a path segment, applying Taylor’s theorem, we can express the energy difference as:
+Where is the angle between the gradient and displacement vectors. For MEPs, the displacement aligns with or against the gradient direction () giving:
 This gives us a surrogate objective with the same minimum:
 Using the triangle inequality on the absolute energy differences, we can establish that minimizing the simple sum of energies along the path provides an upper bound on this action:
+Where is the lowest energy along the path.
 Since the latter terms are constants during optimization, this leads to our final loss function.
@@
 where are uniformly or randomly sampled points in .
-A complete derivation with all mathematical details is provided in Appendix A.1.
 By minimizing , we drive the path representation to discover low-energy regions connecting stable states, approximating the MEP predicted by the Onsager-Machlup framework.
-2.1.1 Neural Representation of Minimum-Energy Paths
+2.1.1 Neural Representation of MEPs
 We propose representing the path using a neural parameterization, similar to the mean parameterization of Du et al.
@@
 While our approaches share similar representations, they differ in objectives and computational efficiency (see Appendix B.1 for theoretical connections and performance comparisons, respectively).
-Specifically, we construct a continuous function mapping a progress coordinate to molecular configurations , where is the number of atoms:
-Here is a neural network, and are basis functions that satisfy the endpoint constraints , , , and , ensuring that and regardless of the network’s output.
+Specifically, we construct a continuous function mapping a progress coordinate to molecular configurations , where is the number of atoms, as illustrated in Figure 1 (left):
+Here is a neural network, and are basis functions that satisfy the endpoint constraints , , , and , ensuring that and regardless of the neural network’s output.
 While simple linear functions and can be used, other basis functions with appropriate boundary conditions are also effective.
-The blending function controls the network’s contribution along the path, typically designed to allow maximum influence in the transition region while smoothly vanishing at endpoints (e.g., ).
+The blending function controls the neural network’s contribution along the path, typically designed to allow maximum influence in the transition region while smoothly vanishing at endpoints (e.g., ).
 2.1.2 Path Optimization
-We minimize directly through stochastic gradient descent. For each batch of points along the path:
-Unlike traditional chain-of-states methods, explicit spring-like continuity terms are unnecessary: the inherent continuity of , combined with smooth blending functions, ensures path continuity without additional constraints.
-2.2 Established Approaches to Minimum-Energy Path Finding
-Classical approaches to minimum-energy path discovery operate directly in Cartesian atomic coordinates, predating differentiable molecular force fields (Wang et al., 2023) that enable our neural network representation.
+We minimize using the Adam optimizer (Kingma & Ba, 2017):
+Unlike traditional chain-of-states methods, explicit spring-like continuity terms are unnecessary; the inherent continuity of , combined with smooth blending functions, ensures path continuity.
+2.2 Established Approaches to MEP Finding
+Classical approaches to MEP discovery operate directly in Cartesian atomic coordinates, predating differentiable molecular force fields (Wang et al., 2023) that enable our neural network representation.
 This direct optimization faces significant challenges as steric clashes create extreme energy barriers that are difficult to overcome with local optimization.
-The string method and chain-of-states approaches discussed below represent broad families with numerous variations, extensively cataloged in reviews such as "Transition-Path Theory and Path-Finding Algorithms for the Study of Rare Events" by E & Vanden-Eijnden (2010).
+The string method and chain-of-states approaches discussed below represent broad families with numerous variations, extensively cataloged in reviews such as by E & Vanden-Eijnden (2010).
 These established techniques illustrate both foundational concepts and limitations that our approach addresses.
 2.2.1 String Method
-The string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Ramakrishnan et al., 2025) determines the MEPs by evolving a curve in the configuration space according to the physical forces while maintaining proper parameterization.
-Given the same endpoints and used in our formulation, the method seeks a curve (string) parameterized by that connects these states.
+The string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007) determines the MEPs by evolving a curve in the configuration space according to the physical forces while maintaining proper parameterization.
+Given the same endpoints and used in our formulation, the method seeks a curve (string) parameterized using splines mapping to the systems Cartesian coordinates that connects these states.
 The MEP satisfies:
@@
 Here is a Lagrange multiplier term enforcing the parameterization constraint, typically equal arc-length parameterization where (constant) for all .
-In practice, the string is discretized into images and evolved through a two-step procedure:
+In practice, the string is discretized into images and evolved through a two-step procedure that requires explicit reparameterization steps:
 1. Evolution step:
@@
 This two-step procedure converges to a discretized MEP satisfying each image’s local orthogonality condition .
-This method requires explicit reparameterization steps that can introduce computational inefficiencies.
 2.2.2 Chain-of-States Methods
-An alternative perspective views MEP finding as optimization of a chain of states (Pratt, 1986; Elber & Karplus, 1987; Ulitsky & Elber, 1990; Jónsson et al., 1998; Henkelman et al., 2000; Gillilan & Lilien, 2004; Sheppard et al., 2008).
+An alternative perspective views MEP finding as optimization of a chain-of-states (Pratt, 1986; Elber & Karplus, 1987; Ulitsky & Elber, 1990; Jónsson et al., 1998; Henkelman et al., 2000; Gillilan & Lilien, 2004; Sheppard et al., 2008).
 Given a sequence with and , the objective function is defined as:
@@
 Although variants of chain-of-states methods exist that avoid this limitation, they come at an increased computational cost.
-The neural representation avoids this limitation through its natural continuity properties.
 3 Experiments
@@
 These systems represent different scales of complexity and serve to validate our method’s ability to discover physically meaningful transition paths.
-In Appendix A.2, we compare and benchmark our method with the related approach outlined in (Du et al., 2024), and in Appendix B.1, we present ablation studies analyzing the impact of different embedding strategies, interpolation methods, network architectures, and regularization approaches.
+In Appendix A.2, we compare and benchmark our method with the related approaches outlined in (Du et al., 2024; Lee et al., 2025).
+In Appendix B.1, we present ablation studies analyzing the impact of different embedding strategies, interpolation methods, neural network architectures, and regularization approaches.
 3.1 AIB9 Transition Path Discovery
@@
 Unlike artificial toy problems, AIB9 allows direct validation with actual biomolecular force fields.
-We use the AMBER ff15ipq-m force field for protein mimetics (Bogetti et al., 2020) implemented in DMFF (Wang et al., 2023), and we generate multiple minimum-energy paths between the system’s two metastable states.
+We use the AMBER ff15ipq-m force field for protein mimetics (Bogetti et al., 2020) implemented in DMFF (Wang et al., 2023), and we generate multiple MEPs between the system’s two metastable states.
 Figure 2 shows the discovered transition paths projected onto the central residue’s - dihedral angle space.
-The method identifies multiple physically plausible pathways between the states, also observed with long MD simulations.
+The method identifies multiple physically plausible pathways between the states, which are also observed with long MD simulations.
 For each path, different conformations of the states were picked as start and end points; this, along with the neural network’s random initialization, leads to the discovery of multiple different transitions.
+Notably, the discovered MEPs align well with the free energy landscape despite our choice of action functional without the trajectory entropy term and the assumption of overdamped dynamics, which may not be ideal for this vacuum system, but is less problematic for hydrated systems.
 4 BPTI Conformational Change Pathway
 Bovine Pancreatic Trypsin Inhibitor (BPTI) is a clinically significant protein used as an anticoagulant in medical procedures.
-BPTI is 58 residues long and has complex conformational changes involving disulfide bond rearrangements paired with backbone changes.
+BPTI is 58 residues long (892 atoms) and has complex conformational changes involving disulfide bond rearrangements paired with backbone changes.
 To investigate BPTI’s conformational transitions, we use as reference structures five key states from a landmark molecular dynamics (MD) simulation performed by D.E.
 Shaw Research on their Anton supercomputer (Shaw et al., 2010).
-These publicly available snapshots provide rare benchmark structures for validating our approach to BPTI’s transition mechanisms.
+These publicly available snapshots provide rare benchmark structures for validating our approach to BPTI’s transition mechanisms, making it an excellent larger system to test the method.
 Full details of our BPTI system preparation, explicit solvent shell construction, and hybrid neural network architecture for handling water molecules are provided in Appendix B.2.
@@
 To analyze this path, we compute two collective variables matching those used in the original D.E.
-Shaw study: (1) the disulfide torsion angle between Cys14 and Cys38, calculated using the dihedral angle formed by CB14–SG14–SG38–CB38 atoms, and (2) the backbone RMSD of residues 4-54 after mean-centering.
-Remarkably, our optimized path passes through all intermediate snapshots, indicating that it successfully discovers the same transition mechanism as the computationally intensive explicit-solvent MD simulation.
+Shaw study: (1) the disulfide torsion angle between Cys14 and Cys38, calculated using the dihedral angle formed by CB14–SG14–SG38–CB38 atoms, and (2) the backbone RMSD, root mean square displacement, of residues 4-54 after mean-centering.
+Remarkably, as shown in Figure 3, our optimized path passes through all intermediate snapshots, indicating that it successfully discovers the same transition mechanism as the computationally intensive MD simulation.
 We matched their simulation conditions by incorporating explicit water molecules in our model to ensure a meaningful comparison with the Anton simulation.
 We generated water shells extending 14Å from any protein atom for both the initial and final BPTI conformations.
-To handle water molecules during transitions, we pair proximal water molecules between start and end states, then model their movement using simple parameterized splines, rather than directly incorporating all water coordinates into the neural network.
+This increases the number of atoms from 892 to 8,300.
+To handle water molecules during transitions, we pair proximal water molecules between start and end states using the Hungarian algorithm (Kuhn, 1955).
+We then model their movement using simple parameterized splines, rather than directly incorporating all water coordinates into the neural network.
 This hybrid neural-spline approach allows us to include explicit solvent effects while keeping computational costs low.
@@
 The system was modeled using the AMBER99sb-ILDN force field (Lindorff-Larsen et al., 2010) with the TIP3P water model (Jorgensen et al., 1983; Neria et al., 1996).
-Each minimum-energy path was computed in approximately 15 minutes on a single A6000 GPU, compared to the D.E.
-Shaw simulation, which required approximately 1.3 million node-hours on their specialized Anton supercomputer to generate the 1.03 millisecond trajectory capturing the same conformational changes.
-While our approach does not provide the same statistical ensemble information as equilibrium MD, it efficiently captures the key transition pathway.
+Each MEP was computed in approximately 15 minutes on a single A6000 GPU, requiring only force field evaluations compared to the D.E.
+Shaw simulation, which conducted approximately 412 billion force field evaluations to generate the 1.03 millisecond trajectory capturing the same conformational changes.
+This represents a six orders of magnitude reduction in computational effort.
+While our approach does not provide the same statistical ensemble information as equilibrium MD, it efficiently captures the key transition pathway, which is often the primary mechanistic insight sought by MD practitioners.
+For most scientifically interesting systems—typically large biomolecular complexes—equilibrium sampling of the complete conformational space or transition path ensemble remains computationally prohibitive, making identification of dominant pathways particularly valuable.
 It’s worth noting that even the reference millisecond-scale simulation may not have fully achieved equilibrium sampling of the conformational space.
@@
 Furthermore, physical force fields introduce challenges through high-energy spikes that can cause training instability and divergence in some runs.
-Our gradient clipping approach mitigates this issue but doesn’t universally prevent training failures.
+Our gradient clipping approach, outlined in B.2.5, largely mitigates this issue but doesn’t universally prevent training failures.
 The method has been validated primarily on small to medium-sized protein systems, and scaling to larger multi-domain proteins could require additional optimization strategies.
-Our approach is also constrained to systems with differentiable force fields, but this limitation diminishes as more implementations become available.
-Finally, while computationally efficient, our method focuses on identifying minimum-energy paths rather than sampling the full statistical ensemble of transition pathways, representing a trade-off between efficiency and exhaustively sampling a system.
+Our approach is also constrained to systems with suitable differentiable force fields, but this limitation diminishes as more implementations become available.
+Finally, while computationally efficient, our method focuses on identifying MEPs rather than sampling the full statistical ensemble of transition pathways, representing a trade-off between efficiency and exhaustively sampling a system.
+However, as noted in the BPTI section 4, this mechanistic insight is often a key objective for MD practitioners studying complex biomolecular transitions.
 6 Conclusion
-We introduced a novel approach to molecular transition path discovery that leverages physics-informed neural networks and differentiable force fields to enable the continuous representation of conformational changes.
-By reformulating minimum-energy path discovery as a neural optimization problem, we demonstrated the method’s effectiveness on both the small AIB9 peptide and the explicitly solvated BPTI protein, showing its ability to discover physically meaningful transition pathways across different complexity scales.
-Our approach achieves remarkable computational efficiency, reducing computational requirements by approximately six orders of magnitude compared to traditional MD simulations for the BPTI system.
-While our method yields a single pathway rather than the complete ensemble of transition trajectories, this represents a valuable trade-off that makes previously intractable conformational transition studies accessible on standard computational resources.
-While our current implementation shows promising results, several directions for improvement remain.
-Future work should explore neural architectures that better respect the structural organization of protein systems.
-Ramakrishnan et al.
-(2025) investigates non-uniform sampling schemes along the path, among other innovative design decisions that might further aid convergence.
+We demonstrated that reformulating MEP discovery as a continuous neural optimization problem enables efficient discovery of molecular transition mechanisms with unprecedented computational efficiency.
+Our approach transforms the complex Onsager-Machlup action functional into a straightforward energy-based loss function that eliminates the need for force and path gradient computations required by related methods, which are known to induce optimization challenges in PINNs.
+This work provides a practical tool for molecular dynamics practitioners that captures essential mechanistic insights without requiring specialized hardware or extensive simulation time.
+While our current implementation shows promising results, several concrete directions for improvement remain.
+Future work should explore neural architectures that incorporate domain-specific structural priors beyond simple MLPs with sinusoidal/Fourier embeddings.
+Related work by Ramakrishnan et al.
+(2025) investigates non-uniform sampling schemes along the path that might further aid convergence for paths that constitute dramatic rearrangements.
 Integrating pretrained generative models suggested by Raja et al.
-(2025) could provide better initial transition guesses, potentially accelerating convergence by starting optimization from more plausible conformational pathways.
-Extending the method to handle larger multi-domain proteins is also a promising frontier.
-Future work could also explore methods to estimate free energies and reaction rates from our discovered paths.
+(2025) could provide better initial transition guesses, potentially accelerating convergence by starting optimization from conformations that already respect molecular geometry.
+Extending the method to handle larger multi-domain proteins is also a frontier of great scientific interest, which might bring new challenges.
+Future work could also explore ways to estimate free energies and reaction rates from our discovered paths.
 7 Broader Impacts
@@
 We introduce stochastic sampling to estimate the continuous energy integral
-For a minimum-energy path, the displacement tends to align either parallel or anti-parallel with the gradient direction (i.e., for path segments)
+For a MEP, the displacement tends to align either parallel or anti-parallel with the gradient direction (i.e., for path segments)
 We use the triangle inequality to establish an upper bound on the sum of absolute energy differences
 While the last two approximations do not change the location of the global minimum, they do affect the optimization landscape before that minimum is reached.
-A.1.1 From Onsager-Machlup to a geometric line integral
+A.1.1 From the Onsager-Machlup Action to a geometric Line Integral
 We begin with a molecular system evolving under overdamped Langevin dynamics, a standard model for high-friction molecular motion:
@@
 with equality when is parallel or anti-parallel to , with the direction determined by whether the path is ascending or descending the energy landscape.
-This is precisely the defining characteristic of the minimum-energy path (MEP) - a path that follows the potential energy landscape’s gradient while avoiding high-energy regions.
+This is precisely the defining characteristic of the MEP - a path that follows the potential energy landscape’s gradient while avoiding high-energy regions.
 When this equality holds, our action simplifies to:
@@
 This represents a transition from the theoretical derivation using discrete segments to a practical implementation using sampled points for more efficient optimization.
-We employ stochastic sampling of the progress variable to construct our loss function:
+We employ stochastic sampling of the progress coordinate to construct our loss function:
 where are uniformly sampled progress coordinates and is the batch size (the number of points sampled along the path during each optimization step).
 This stochastic approach provides an unbiased estimator of the expected energy along the path:
-By minimizing through gradient-based optimization, the neural network learns to represent low-energy pathways connecting the stable states, effectively discovering the minimum-energy path predicted by the Onsager-Machlup framework.
+By minimizing through gradient-based optimization, the neural network learns to represent low-energy pathways connecting the stable states, effectively discovering the MEP predicted by the Onsager-Machlup framework.
 The resulting path follows the potential energy landscape’s gradient while avoiding high-energy regions, representing the most probable transition mechanism in the overdamped limit.
-A.2 Connection between Doob’s Lagrangian and Minimum-Energy Path Optimization
+A.2 Connection between Doob’s Lagrangian and MEP Optimization
 We elaborate on the theoretical connections between our energy-based optimization approach and the framework of Doob’s h-transform for sampling transition paths (Du et al., 2024).
@@
 We provide a direct comparison between our PINN-MEP approach and the method of Du et al.
-in terms of computational efficiency, as evaluated by the number of force-field/energy function calls, and the maximum energies observed in the paths, corresponding to the peak of the potential barrier.
+in terms of computational efficiency, as evaluated by the number of force field/energy function calls, and the maximum energies observed in the paths, corresponding to the peak of the potential barrier.
 To provide a rigorous comparison, we conducted benchmarks using experimental conditions identical to those reported by Du et al.
@@
 Appendix B Comparison with other related Methods
-While our primary benchmark comparison focuses on the Doob’s Lagrangian method due to similarity in task and system setup, examining how our approach relates to other methods for transition path discovery/minimum energy path generation is also informative.
-Here we present results from two related approaches in the literature, though we emphasize that direct numerical comparisons should be treated with caution due to differences in force fields, optimization objectives, and computational frameworks.
-| System | Method | Total energy evals | TS energy (eV) |
-| Alanine dipeptide | NEB | 8063 | 11.0014 |
+While our primary benchmark comparison focuses on the Doob’s Lagrangian method due to similarity in task and system setup, examining how our approach relates to other methods for transition path discovery/MEP generation is also informative.
+Here we present results from two related approaches in the literature, though we emphasize that direct numerical comparisons should be treated with caution due to differences in force fields, optimization objectives, and units used.
+Method | # Evaluations | Max Energy (eV) |
+NEB | 8063 | 11.0014 |
 INR | 3046 | 0.4070 |
-| Method | CVs | # FF Evals / Path | Runtime / Path |
+Method | CVs | # Evaluations | Runtime / Path |
 MCMC (Two-Way Shooting) | No | 1B | 100 hours |
@@
 Table 2 shows results from Ramakrishnan et al.
-(2025) comparing Nudged Elastic Band (NEB) with their Implicit Neural Representation (INR) approach for minimum-energy path discovery.
+(2025) comparing Nudged Elastic Band (NEB) with their Implicit Neural Representation (INR) approach for MEP discovery.
 Both our PINN-MEP approach and their INR method utilize lightweight, preexisting molecular dynamics force fields, though we employ different such force fields.
-Their approach requires a more complex loss function that evaluates both energy and forces along with path parameterization and its gradient (path velocity).
+Their approach requires a more complex loss function that evaluates both energy and forces along with path parameterization and its gradient (path velocity), but needs fewer force field evaluations than our method.
+It is important to note that the "Max Energy" reported in Table 2 is measured in electron volts (eV) and appears to be adjusted for ground state energy reference, which makes direct energy comparisons with our method challenging.
+The difference in force fields used further complicates any direct energy value comparisons between methods.
 In contrast, Table 3 presents results from Raja et al.
-(2025), who take a fundamentally different approach by repurposing pre-trained generative models to perform sampling of transition paths.
+(2025), who take a fundamentally different approach by re-purposing pre-trained generative models to perform sampling of transition paths.
 A key advantage of their method is that the generative model provides significantly better initial guesses for transition paths based on learned molecular distributions.
-However, their approach requires evaluating both the pretrained model (as a surrogate force field) and its gradient for their loss function.
-Our PINN-MEP method stands out for its simplicity, using only the energy potential in the loss function without requiring additional gradient terms, while maintaining comparable computational efficiency.
-These tables corroborate our findings in the Doob’s Lagrangian comparison: methods optimizing for a single minimum-energy path consistently require orders of magnitude fewer force field evaluations than approaches that characterize the full transition path ensemble.
+Their approach requires evaluating both the pretrained model (as a surrogate force field) and its gradient for their loss function, which is more computationally intense due to the high number of parameters of the pretrained model as compared to a physics-derived force field.
+These tables corroborate our findings in the Doob’s Lagrangian comparison: methods optimizing for a single MEP consistently require orders of magnitude fewer force field evaluations than approaches that characterize the full transition path ensemble.
 B.1 Ablation Studies
@@
 We conduct these ablation studies on the AIB9 system because Alanine dipeptide, as a benchmark system, is too simple to provide meaningful insights into the performance of our method.
-For each ablation study, we report the lowest maximum energy of the path, the mean and standard deviation of the energy, and the mean number of force field evaluations to reach the lowest maximum energy.
-We run each setting in the ablations five times for a maximum of 150000 batches, with a batch size of 1 and 32 frames per batch.
-For experiments without a spring constant between adjacent sampled frames, the batch size and frames are treated identically.
+For each ablation study, we report the lowest maximum energy of the path among all training runs, the mean and standard deviation of the energy, and the mean number of iterations to reach the lowest maximum energy.
+We run each setting in the ablations 15 times for a maximum of 25,000 iterations.
+We use a batch size of 1, where each batch contains 32 values of the progress coordinate for most cases.
+The exception is when we evaluate the chain-of-states-like regularization, which minimizes the difference between network outputs for adjacent values of ; in this case, the effective batch size is 32, since there are no interactions between different network outputs otherwise.
 Due to the nature of physical force fields, slight deviations in the path can lead to large changes in energy, leading to similar settings having very different energies.
@@
 However, we find that the results are still directionally informative.
-B.1.1 Embedding Strategy
-We compared sinusoidal time embeddings with different maximum periods against simple linear projections of the progress variable.
-Table 4 shows sinusoidal embeddings with a period of generally outperformed other configurations.
-| Embedding Type | Max Period | Mean Energy | Best Energy | Evaluations |
-| Sinusoidal | 5832.29 7006.09 | -33.80 | 142400 | |
-| Sinusoidal | 4939.20 2836.56 | 1388.48 | 122100 | |
-| Sinusoidal | 9508.61 7297.49 | 1273.18 | 106900 | |
-| Sinusoidal | 5851.25 4618.35 | 819.25 | 129600 | |
-| Linear (None) | – | 9567.01 7578.20 | 303.63 | 127600 |
-B.1.2 Spring Regularization
+B.1.1 General Settings for Ablation Studies
+All ablation studies are conducted using the following base configuration, with each study varying only specific parameters while keeping all others fixed:
+Molecular System Parameters:
+Molecular system: AIB9 peptide with 129 atoms
+Force field: AMBER ff15ipq-m for protein mimetics
+Cutoff method: No cutoff applied
+Simulation box size: 10.0 nm
+Neighbor list cutoff distance: 4.0 nm
+Training Configuration:
+Number of iterations: 25,000 per run
+Number of independent runs: 15 per parameter setting
+Evaluation frequency: Every 100 training steps
+Optimizer: Adam with optax default parameters
+Learning rate:
+Path Sampling Parameters:
+Batch size: 1
+Frames per batch: 32 points sampled along the progress coordinate
+Spring constant: (no chain-of-states regularization)
+Interpolation method: Cubic Hermite for blending functions
+Network Architecture (Base Configuration):
+Network depth: 6 hidden layers
+Width factor: 4 (network width scales as number of atoms)
+Embedding dimension: Automatically set to
+Time embedding: Sinusoidal with maximum period of
+Activation function: ReLU throughout the network
+Weight initialization: Glorot uniform for hidden layers, zero initialization for final layer
+Bias initialization: Default for hidden layers, zero initialization for final layer
+B.1.2 Embedding Strategy
+We compared sinusoidal time embeddings with different maximum periods against simple linear projections of the progress coordinate.
+Table 4 shows that linear embedding (None) performed best, achieving the lowest mean energy, though sinusoidal embeddings with period also achieved competitive results.
+We find these results at odds with the need for sinusoidal embeddings in the larger BPTI system.
+Perhaps the AIB9 system, which is characterized by one large general conformational rearrangement primarily driven by main-chain dihedral angles, does not need the high frequency features that sinusoidal embeddings provide, but which might be essential for other systems, due to the well-known bias of MLPs toward modeling low frequencies in data.
+Embedding | Max Period | MeanMax Energy () | MinMax Energy | Mean Iterations |
+Linear (None) | – | 591.35 963.72 | -337.36 | 200873 |
+Sinusoidal | 1326.13 2339.35 | -337.36 | 210907 | |
+Sinusoidal | 1528.18 2429.05 | -129.90 | 226290 | |
+Sinusoidal | 1888.66 3604.47 | -337.36 | 187493 | |
+Sinusoidal | 2453.53 3507.47 | -337.36 | 217053 |
+B.1.3 Spring Regularization
 We tested the effect of chain-of-states-like spring constant regularization, as shown in Table 5.
-Interestingly, we found that a small amount of spring regularization () provided the best results, suggesting that a minimal amount of explicit path continuity enforcement can be beneficial when combined with our neural representation, however, this is not a requirement for the method to work and the additional hyperparameter does not necessarily be worth including.
-| Spring Constant | Mean Energy | Best Energy | Evaluations |
-| 368.78 416.68 | -278.09 | 130300 | |
-| 8862.90 9138.59 | -155.85 | 120600 | |
-| 0.0 | 8200.08 15576.30 | 240.85 | 116500 |
-| 1156.70 428.50 | 791.59 | 120400 | |
-| 7440.13 3339.87 | 4657.28 | 92300 | |
-| 45567.30 5015.78 | 40609.10 | 64700 |
-B.1.3 Network Architecture
+Interestingly, we found that small amounts of spring regularization ( or ) provided the best results, with achieving the lowest mean energy overall.
+This suggests that a minimal amount of explicit path continuity enforcement can be beneficial when combined with our neural representation, though this is not a strict requirement for the method to work.
+However, adding a spring constant reintroduces intra-batch dependencies, which means that batch sizes should be of a minimum size for this approach to work effectively.
+With low batch sizes, this could lead to the corner cutting effects commonly observed in traditional chain-of-states methods, which might make scaling the addition of spring regularization difficult for larger systems.
+Since we are able to reach the same minimum energy without spring regularization, we recommend against its use, especially when scaling to larger systems where lower batch sizes are preferred.
+Spring Constant | MeanMax Energy () | MinMax Energy | Mean Iterations |
+-178.22 260.09 | -337.36 | 153533 | |
+-155.41 286.63 | -337.36 | 101600 | |
+803.61 586.95 | 205.74 | 159033 | |
+0.0 | 1460.41 2486.05 | -337.36 | 209133 |
+10528.4 7394.89 | 5065.46 | 155400 | |
+48639.7 12111.7 | 33143.6 | 107267 |
+B.1.4 Network Architecture
 We investigated the effect of network depth and width on transition path quality.
-Table 6 shows that an 8-layer network with a width factor of 4 achieved the best balance between expressivity and computational efficiency, with wider networks also performing well.
-| Depth | Width Factor | Mean Energy | Best Energy | Evaluations |
-| 8 | 4 | 2746.08 5381.57 | -168.17 | 112300 |
-| 8 | 6 | 6337.10 6284.13 | -278.09 | 79300 |
-| 8 | 2 | 10396.20 14039.40 | -278.09 | 113500 |
-| 6 | 6 | 6189.19 7667.11 | 334.22 | 112800 |
-| 6 | 4 | 9371.15 15277.50 | 240.85 | 111800 |
-| 6 | 2 | 9872.01 14229.30 | 152.36 | 122500 |
-| 4 | 6 | 13242.00 8432.26 | 6013.58 | 94000 |
-| 4 | 4 | 16300.90 5768.37 | 8295.52 | 132100 |
-| 4 | 2 | 24916.70 14933.30 | 12697.60 | 78700 |
-B.1.4 AIB9 Network Architecture
+Table 6 shows that an 8-layer network with a width factor of 4 achieved the best mean energy performance, followed closely by a 6-layer network with width factor 6.
+Depth | Width Factor | MeanMax Energy () | MinMax Energy | Mean Iterations |
+8 | 4 | 25.02 467.23 | -337.36 | 159333 |
+6 | 6 | 72.42 458.72 | -337.36 | 210267 |
+8 | 6 | 165.91 1021.79 | -337.36 | 219433 |
+8 | 2 | 1434.97 3206.06 | -337.36 | 178167 |
+6 | 4 | 1460.34 2486.07 | -337.36 | 209067 |
+6 | 2 | 3621.8 3198.25 | 217.83 | 231200 |
+4 | 6 | 4242.74 3568.51 | 140.81 | 231967 |
+4 | 4 | 5130.75 5758.01 | -328.49 | 228133 |
+4 | 2 | 16200.5 8032.18 | 159.21 | 210367 |
+B.1.5 Interpolation Type
+We examined different interpolation methods for the blending functions that combine the neural network output with the endpoint constraints.
+Table 7 compares linear, cosine, quintic, and cubic Hermite interpolation.
+Cosine interpolation performed best in terms of mean energy, though all methods were able to achieve the minimum energy of -337.36 in at least one run, except for quintic.
+The cubic Hermite interpolation showed higher variance in results.
+The specific formulas for each interpolation type are as follows:
+Linear interpolation:
+Cubic Hermite interpolation:
+Quintic interpolation:
+Cosine interpolation:
+where and are the basis functions that ensure endpoint constraints and , and is the blending function that controls the neural network’s contribution along the path.
+Interpolation Type | MeanMax Energy () | MinMax Energy | Mean Iterations |
+Cosine | 442.51 769.51 | -337.36 | 198720 |
+Linear | 602.61 1057.13 | -337.36 | 219147 |
+Quintic | 967.16 1288.57 | -281.14 | 210227 |
+Cubic Hermite | 1325.49 2339.75 | -337.36 | 212807 |
+B.1.6 AIB9 Network Architecture
+Below are the main details of the AIB9 MEP training runs, as shown in the main paper.
 Model type: Transition MLP with sinusoidal time embedding
@@
 Activation function: ReLU
-B.1.5 Training Parameters
+B.1.7 Training Parameters
 Optimizer: Adam
-Learning rate: 0.00001
+Learning rate:
 Batch size: 4
@@
 Interpolation method: Linear
+All other parameters are identical to the base settings listed in the ablations section B.1.1.
 B.2 BPTI Modeling and Computational Details
@@
 Shaw Research BPTI simulation of BPTI (Shaw et al., 2010).
-The initial and final states for our minimum-energy path optimization were selected as the second snapshot (index 1) and the second-to-last snapshot (index -2) from the provided reference structures, as they were the most separated in the CVs determined in the original study, named the backbone RMSD and the disulfide torsion angle.
+The initial and final states for our MEP optimization were selected as the second snapshot (index 1) and the second-to-last snapshot (index -2) from the provided reference structures, as they were the most separated in the CVs determined in the original study, named the backbone RMSD and the disulfide torsion angle.
 B.2.2 Explicit Solvent Shell Construction
@@
 Total iterations: 30,000
+All other parameters are identical to the base settings listed in the ablations section B.1.1.
 B.2.5 Handling High-Energy Spikes During Training
@@
 Shaw Research simulation on their specialized Anton supercomputer, which required approximately 1.3 million node-hours to generate the 1.03 millisecond trajectory capturing the same conformational changes.
+B.2.8 Computational Resources
+All experiments reported in this paper were conducted on a single NVIDIA A6000 GPU with 48GB of memory.
+Additional computational resources were used during the development phase of this project for hyperparameter tuning across network architectures and optimization parameters.

### v6

@@
-PINN-MEP: Continuous Neural Representations for Minimum-Energy Path Discovery in Molecular Systems
+Follow the MEP: Scalable Neural Representations for Minimum-Energy Path Discovery in Molecular Systems
 Abstract
-Characterizing conformational transitions in physical systems remains a fundamental challenge, as traditional sampling methods struggle with the high-dimensional nature of molecular systems and high energy barriers between stable states.
-These rare events often represent the most biologically significant processes, like ion channel opening that controls cellular signaling, yet may require months of continuous simulation to observe.
-We present a method that reformulates transition path generation as a continuous optimization problem solved through physics-informed neural networks (PINNs) inspired by string methods for minimum-energy path (MEP) generation.
-By representing paths as implicit neural functions and leveraging automatic differentiation with differentiable molecular force fields, our method efficiently discovers physically realistic transition pathways without expensive sampling.
-We demonstrate our approach on two proteins, including an explicitly hydrated BPTI system with more than 8,300 atoms.
-Our method identifies a representative minimum-energy path for the conformational change captured in a millisecond-scale MD simulation (Shaw et al., 2010), while requiring only force field evaluations compared to approximately 412 billion in the original study.
-Though our method yields a single pathway rather than an ensemble of transition trajectories, this represents a six orders of magnitude reduction in computational cost, generating a physically meaningful transition in less than 15 minutes on a standard GPU rather than weeks on specialized hardware.
+Characterizing conformational transitions in physical systems remains a fundamental challenge, as traditional sampling methods struggle with the high-dimensional nature of molecular systems and high-energy barriers between stable states.
+These rare events often represent the most biologically significant processes, yet may require months of continuous simulation to observe.
+One way to understand the function and mechanics of such systems is through the minimum energy path (MEP), which represents the most probable transition pathway between stable states in the high-friction, low-temperature limit.
+We present a method that reformulates MEP discovery as a fast and scalable neural optimization problem.
+By representing paths as implicit neural representations and training with differentiable molecular force fields, our method discovers transition pathways without expensive sampling.
+Our approach scales to large biomolecular systems through a simple loss function derived from the path’s likelihood via the Onsager-Machlup action and a scalable new architecture, AdaPath.
+We demonstrate this approach on two proteins, including an explicitly hydrated BPTI system with more than 3,500 atoms.
+Our method identifies a MEP that captures the same conformational change observed in a millisecond-scale molecular dynamics (MD) simulation in just minutes on a standard GPU, rather than weeks on a specialized cluster.
 1 Introduction
-Characterizing molecular conformational changes, or transitions, is essential for understanding biological processes, from enzyme catalysis to protein folding.
-Although molecular dynamics (MD) simulations provide unprecedented atomic-level detail of these processes (Dror et al., 2012), capturing rare transition events remains computationally challenging due to the inherent timescale gap between molecular motion and conformational changes.
+When studying the function of large biological systems, molecular dynamics simulations effectively serve as a computational microscope, providing atomic-level detail and control that are impossible with experimental methods alone (Dror et al., 2012).
+Their true value often lies in not exhaustive sampling of the system’s Boltzmann distribution—which is intractable for such systems—but in generating hypotheses from single observed transitions that can then be corroborated through targeted in vitro or in vivo experiments.
+However, capturing even these individual transition events remains computationally challenging due to the inherent timescale gap between molecular motion and conformational changes.
 This difficulty arises from transition rates between stable states being exponentially suppressed by the height of the energy barriers that separate them, as described by transition state theory (Eyring, 1935).
-As a result, direct simulation methods spend the vast majority of computational effort sampling already known stable states, making the observation of transitions between them prohibitively expensive.
+As a result, direct simulation methods spend the vast majority of computational effort sampling already known stable states, making the observation of transitions between them prohibitively expensive, often requiring months of continuous simulation to observe a single biologically significant event.
 The search for efficient methods to identify transition pathways between stable states has a rich history in computational physics.
 Early theoretical work established key principles through the Freidlin-Wentzell (Freidlin & Wentzell, 1998) and Onsager-Machlup (Onsager & Machlup, 1953) functionals, which connect the most probable trajectories, including transitions, to minimizers of specific action functionals.
-Building on these foundations, the minimum-energy path (MEP) between stable molecular states represents the most probable transition pathway in the limit of high friction.
-This path reveals the mechanism of conformational change, identifies key transition states and intermediate configurations, and can thus be highly informative about the system’s dynamics.
-Moreover, the MEP can serve as a scaffold for subsequent sampling of the full transition path ensemble, making it a powerful tool for understanding molecular dynamics.
-Methods such as the string method family or chain-of-states approaches (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Petersen et al., 2024; Dellago et al., 1998) approximate these paths as splines or as a discrete series of molecular configurations connected by artificial spring forces to ensure continuity.
-Despite their strong theoretical foundation, these approaches face fundamental challenges stemming from their optimization in Cartesian molecular coordinates.
-The discrete chain of conformations requires periodic reparameterization to maintain even spacing, leading to a complex alternation between energy minimization and geometric constraints.
-Additional limitations include the difficulty of generating good initial transition guesses and limited parallelization, which restricts the scale of practical applications.
+Building on these foundations, the minimum-energy path (MEP) between stable molecular states represents the most probable transition pathway in the limit of high friction and low temperature.
+This path reveals the mechanism of conformational change, identifies key transition states and intermediate configurations, and can thus be highly informative about a system’s dynamics.
+Moreover, the MEP can serve as a scaffold for subsequent sampling of the whole transition path ensemble, making it a powerful tool for understanding molecular dynamics.
+Methods such as the string method family or chain-of-states approaches (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Petersen et al., 2024; Dellago et al., 1998) approximate these paths as a discrete series of molecular configurations connected either by artificial spring forces or held together by frequent reparametrization during optimization to ensure continuity.
+Despite their strong theoretical foundation, these approaches face fundamental challenges stemming from their optimization in Cartesian molecular coordinates (Ovchinnikov et al., 2011).
+The discrete chain of conformations requires either slow periodic reparameterization or a spring force that introduces a hyperparameter to maintain even spacing (Lindgren et al., 2019).
+Additional limitations include the difficulty of generating good initial transition guesses (Ovchinnikov et al., 2011) and limited parallelization, which restricts the scale of practical applications.
 While this formulation has proven useful for small-scale systems, it has not yet benefited from modern optimization approaches that could address these challenges.
-Recent advances in machine learning methodology have created opportunities for novel approaches to this challenge.
-Physics-informed neural networks (PINNs) have increased our ability to solve complex physical problems by embedding differential equations directly into neural architectures (Raissi et al., 2019).
-Additionally, the development of differentiable molecular force fields (Wang et al., 2023) has enabled seamless integration of physical force field models used in MD simulations with modern machine learning frameworks, allowing efficient computation of derivatives beyond forces, which are needed for neural path optimization.
+Recent advances in machine learning have opened the door to strategies addressing this problem.
+Physics-informed neural networks (PINNs) have shown that neural networks can be trained directly on physical laws expressed as differential equations (Raissi et al., 2019).
+In addition, the emergence of differentiable molecular dynamics force fields (Wang et al., 2023) has extended this principle to molecular simulations.
+Together, these developments provide a foundation for deep learning or other optimization approaches that require efficient computation of derivatives beyond forces.
 Numerous approaches have emerged over the past decades to tackle the problem of sampling transitions.
-Traditional simulation-based sampling methods (Jung et al., 2023; Lazzeri et al., 2023; Bolhuis et al., 2002; Dellago et al., 1998; Laio & Parrinello, 2002; Torrie & Valleau, 1977) rely on expensive molecular dynamics simulations, often requiring initial MD-harvested transition trajectories or prior knowledge of the transition dynamics to begin with.
+Traditional simulation-based sampling methods (Jung et al., 2023; Lazzeri et al., 2023; Bolhuis et al., 2002; Dellago et al., 1998; Laio & Parrinello, 2002; Torrie & Valleau, 1977) rely on expensive molecular dynamics simulations, often requiring initial MD-harvested transition trajectories or prior knowledge of the transition dynamics.
 More recent machine learning approaches have explored several directions: variational approaches using Doob’s h-transform (Du et al., 2024; Lee et al., 2025; Das & Limmer, 2019; Singh & Limmer, 2023), stochastic control formulations (Yan et al., 2022; Holdijk et al., 2023), and even attempts to sample transition trajectories directly using diffusion models (Petersen et al., 2023; Han et al., 2024; Jing et al., 2024).
 Notably, Du et al.
-(2024) uses a similar neural representation of transition paths but with a different objective, targeting the complete transition path ensemble by additionally parameterizing path width and optimizing based on a framework rooted in Doob’s h-transform (Doob, 1957) and Wasserstein Lagrangian flows (Neklyudov et al., 2024).
+(2024) also uses a neural representation of transition paths but with a different objective, targeting the complete transition path ensemble by additionally parameterizing path width and optimizing based on a framework rooted in Doob’s h-transform (Doob, 1957) and Wasserstein Lagrangian flows (Neklyudov et al., 2024).
 Most closely related to our method, Ramakrishnan et al.
-(2025) employ a continuous neural parameterization of MEPs inspired by Nudged Elastic Band (NEB) methods, a class of methods related to string methods, incorporating specialized components for tangential forces, path velocity, and transition state energetics to optimize the neural path.
-Another promising concurrent approach by Raja et al.
-(2025) creatively leverages pretrained diffusion models as a sampler of initial transition conformations and as a force field to generate plausible transition paths using the discretized Onsager-Machlup action directly as a loss function.
-While each approach offers valuable insights into neural network-based path discovery, our method focuses on a streamlined optimization scheme that efficiently scales to much larger biomolecular systems with explicit solvent, essential for accurately modeling such systems, while maintaining fast training times.
-Drawing inspiration from neural implicit representations in computer vision (Mildenhall et al., 2020; Sitzmann et al., 2020), we reformulate the traditional minimum-energy-path finding problem as a continuous optimization framework.
-Our approach transforms the Onsager-Machlup functional, which naturally models the probability of paths in stochastic systems, into a remarkably simple and intuitive loss function that requires only energy evaluations along the neural network parametrized path.
-In contrast to existing methods that use complex composite losses involving energies, forces, and path gradients, our streamlined formulation eliminates the need for these additional computations.
-This simplification, while faster, also sidesteps known issues in PINN optimization if the loss contains differential operators (Rathore et al., 2024).
-This PINN framework efficiently captures physically meaningful transition mechanisms for complex biomolecular systems, completing optimization in minutes on standard hardware.
+(2025) employs a continuous neural parameterization of MEPs inspired by Nudged Elastic Band (NEB) methods, incorporating specialized components for tangential forces, path velocity, and transition state energetics.
+Another concurrent approach by Raja et al.
+(2025) leverages pretrained diffusion models as both samplers and force fields.
+However, these neural approaches face two key limitations: they employ standard MLP architectures that scale poorly with system size or large pretrained models, and they rely on complex multi-term loss functions involving higher-order differential terms known to hamper optimization (Rathore et al., 2024).
+Drawing inspiration from neural implicit representations in computer vision (Mildenhall et al., 2020; Sitzmann et al., 2020), we aim to address both limitations.
+We derive a simple, single-term energy-based loss directly from the Onsager-Machlup functional and introduce AdaPath, a novel architecture that employs parameter sharing across atoms.
+Unlike existing approaches that enforce continuity through complex loss terms, our neural representation achieves smooth paths inherently through architectural design and optimizer choice.
+This combination enables efficient scaling to large biomolecular systems while maintaining fast, stable optimization.
+We further perform extensive multi-system benchmarks of our method against classical and deep learning approaches, and find that our approach achieves far lower path and transition state energies.
 2 Theory and Methods
 2.1 Derivation of the Energy-Based Loss Function and Neural Implementation
-To derive our loss function, which when minimized yields a continuous curve representing the MEP connecting fixed endpoints and , we build upon alternative formulations of the Onsager-Machlup action functional as explored by Vanden-Eijnden & Heymann (2008) and Olender & Elber (1997) in the context of string methods.
+To derive our loss function, which, when minimized, yields a continuous curve representing the MEP connecting fixed endpoints and , we build upon alternative formulations of the Onsager-Machlup action functional as explored by Vanden-Eijnden & Heymann (2008) and Olender & Elber (1997) in the context of string methods.
 However, we further simplify the approach after reaching their geometric formulation.
 A complete derivation with all mathematical details is provided in Appendix A.1.
-We begin with a molecular system evolving under overdamped Langevin dynamics, a model for molecular motion in the high-friction regime characteristic of biomolecular systems:
+To introduce the Onsager-Machlup action intuitively, we begin with a molecular system evolving under overdamped Langevin dynamics, a model for molecular motion in the high-friction regime characteristic of biomolecular systems:
 represents the potential energy function, is Gaussian white noise modeling thermal fluctuations, and physical constants are absorbed into the nondimensionalization.
@@
 However, the resultant MEPs still empirically align well with the free energy landscape.
+Notably, this might not be the case for strongly entropy-driven systems such as intrinsically disordered proteins, among others (Granata et al., 2015; Caro et al., 2017; Skriver et al., 2023).
 To further simplify, we expand the integrand of the action functional:
 For paths with fixed endpoints, the cross-term integrates to a constant difference in potential energy:
-Since and are fixed in our formulation, this term becomes a constant that doesn’t affect the optimization.
+Since and are fixed in our formulation, this term becomes a constant that does not affect the optimization.
 We can therefore focus on minimizing the remaining terms.
@@
 This geometric integral represents a natural endpoint for theoretical development for traditional chain-of-states and string methods.
-In these classical approaches, forces directly update molecular conformations to minimize , making minimization along a discretized string using forces conceptually straightforward.
+In these classical approaches, forces directly update molecular conformations to minimize , making minimization along a discretized string conceptually straightforward.
 However, we want a loss function that acts directly on the neural network parameters in our neural network formulation.
@@
 We represent the transition path as a continuous neural mapping from a progress coordinate to the system’s Cartesian coordinates, such that and . We then discretize the path into segments:
-where and .
-For a path segment, applying Taylor’s theorem, we can express the energy difference as:
-Where is the angle between the gradient and displacement vectors. For MEPs, the displacement aligns with or against the gradient direction () giving:
-This gives us a surrogate objective with the same minimum:
+where and . For a path segment, applying Taylor’s theorem, we can express the energy difference as:
+Where is the angle between the gradient and displacement vectors.
+For MEPs, the displacement aligns with or against the gradient direction ().
+Critically, this approximation requires to be sufficiently small for the higher-order term to be negligible.
+This yields:
+This gives us a surrogate objective with the same minimum, provided the Taylor series approximation holds:
 Using the triangle inequality on the absolute energy differences, we can establish that minimizing the simple sum of energies along the path provides an upper bound on this action:
@@
 To implement this in practice, we sample B points along a parametrized path and directly minimize the sum of energies at these points, yielding:
-where are uniformly or randomly sampled points in .
-By minimizing , we drive the path representation to discover low-energy regions connecting stable states, approximating the MEP predicted by the Onsager-Machlup framework.
-2.1.1 Neural Representation of MEPs
-We propose representing the path using a neural parameterization, similar to the mean parameterization of Du et al.
-(Du et al., 2024).
-While our approaches share similar representations, they differ in objectives and computational efficiency (see Appendix B.1 for theoretical connections and performance comparisons, respectively).
-Specifically, we construct a continuous function mapping a progress coordinate to molecular configurations , where is the number of atoms, as illustrated in Figure 1 (left):
-Here is a neural network, and are basis functions that satisfy the endpoint constraints , , , and , ensuring that and regardless of the neural network’s output.
-While simple linear functions and can be used, other basis functions with appropriate boundary conditions are also effective.
-The blending function controls the neural network’s contribution along the path, typically designed to allow maximum influence in the transition region while smoothly vanishing at endpoints (e.g., ).
-2.1.2 Path Optimization
-We minimize using the Adam optimizer (Kingma & Ba, 2017):
-Unlike traditional chain-of-states methods, explicit spring-like continuity terms are unnecessary; the inherent continuity of , combined with smooth blending functions, ensures path continuity.
+where are uniformly sampled points in .
+While this energy-based loss successfully identifies MEPs when the underlying assumptions hold, it admits a pathological minimum when they break down.
+Specifically, if is large, the validity of the Taylor series approximation breaks down.
+In that case, the loss can be minimized by paths that remain in one stable state and then rapidly ”teleport” to the other, avoiding the high-energy transition region.
+To prevent this pathological solution, we employ specific architectural choices and optimizer selection to control that we detail in the following section.
+2.1.1 AdaPath: A Scalable Neural Architecture for MEPs
+Neural approaches to modeling transitions map directly to , requiring hidden layers that scale with system dimensionality, leading to quadratic parameter scaling that limits applicability to large systems.
+AdaPath 1 addresses this through parameter sharing across atoms.
+Each atom has a learnable embedding , processed through shared transformer-like MLP blocks with adaptive layer norm (adaLN) conditioning (Peebles & Xie, 2023).
+A conditioning network processes sinusoidal embeddings of to generate scale , shift , and gate parameters that modulate each MLP block , and in addition to a final scale and shift before the final projection layer that projects to the atoms’ Cartesian coordinates .
+Neural implicit representations (NIR) effectively function as smooth lookup tables with a fixed range of inputs rather than complex mappings as in conventional deep learning tasks, allowing neural networks to remain small while embeddings drive the parameter count.
+In AdaPath, this enables conditioning parameters, which are computed once per value, to be heavily parameterized, while keeping the main branch operating on each atom embedding lightweight.
+AdaPath maintains endpoint constraints through:
+where and provide linear interpolation ensuring and , and controls network contribution, vanishing at endpoints.
+The residual MLP design maintains path continuity and enables gradient flow.
+While our energy-based loss does not explicitly enforce continuity, residual connections preserve smooth transitions by allowing direct information flow from embeddings to output.
+However, standard optimizers can cause excessive path stretching around transition regions, see 6.
+We therefore use the Muon optimizer for small systems (Keller Jordan et al., 2024), as it maintains lower Lipschitz constants (Large et al., 2024), ensuring adequate path presence in transition regions for accurate energy estimates.
+For large systems, Adam (Kingma & Ba, 2017) suffices; we posit that this might be due to neural networks’ increasing inability to model discontinuities as dimensionality increases (Petersen & Voigtlaender, 2018).
+We further scale the energy function and use gradient clipping, as outlined in H, to avoid high energy spikes caused by steric clashes between atoms.
 2.2 Established Approaches to MEP Finding
-Classical approaches to MEP discovery operate directly in Cartesian atomic coordinates, predating differentiable molecular force fields (Wang et al., 2023) that enable our neural network representation.
-This direct optimization faces significant challenges as steric clashes create extreme energy barriers that are difficult to overcome with local optimization.
-The string method and chain-of-states approaches discussed below represent broad families with numerous variations, extensively cataloged in reviews such as by E & Vanden-Eijnden (2010).
-These established techniques illustrate both foundational concepts and limitations that our approach addresses.
-2.2.1 String Method
-The string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007) determines the MEPs by evolving a curve in the configuration space according to the physical forces while maintaining proper parameterization.
-Given the same endpoints and used in our formulation, the method seeks a curve (string) parameterized using splines mapping to the systems Cartesian coordinates that connects these states.
-The MEP satisfies:
-where is the component of the force perpendicular to the unit tangent vector along the string.
+All MEP discovery methods must solve a fundamental optimization challenge: finding paths that simultaneously minimize energy while maintaining smooth continuity between endpoint configurations.
+This dual objective creates an inherent tension; energy minimization alone may produce discontinuous jumps between low-energy regions, while continuity enforcement alone may force paths through high-energy barriers.
+Different methodological families resolve this through distinct mathematical formulations, each with characteristic advantages and limitations that our neural approach addresses.
+The approaches discussed below represent broad families with numerous variations, extensively cataloged in reviews such as E & Vanden-Eijnden (2010).
+2.2.1 Explicit Continuity Penalties
+The chain-of-states family (Pratt, 1986; Elber & Karplus, 1987; Ulitsky & Elber, 1990) handles the energy-continuity trade-off through composite objective functions that explicitly balance both terms.
+Given a sequence with and , the optimization targets:
+Where the first term enforces energy minimization and the spring term constant penalizes discontinuities.
+However, this formulation suffers from ”corner cutting”, where the path takes shortcuts through high-energy regions when spring forces dominate.
+The Nudged Elastic Band (NEB) method (Jónsson et al., 1998; Henkelman et al., 2000) refines this approach by projecting spring forces onto the path normal direction, yielding:
+Where represents forces perpendicular to the path and the spring term acts only parallel to it.
+This decoupling reduces corner cutting but requires careful force projection and still faces scaling challenges in high-dimensional systems, often leading to jagged paths (Zhang et al., 2016).
+2.2.2 Continuity via Geometric Constraints
+The string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006) takes an alternative approach, separating energy minimization from continuity enforcement through alternating optimization.
+The method seeks a curve satisfying the orthogonality condition:
+where is the unit tangent vector along the string.
 This orthogonality condition has an intuitive interpretation: at each point along the MEP, the forces acting on the system are entirely parallel to the path itself, with no perpendicular components trying to push the system away from the path.
-This represents a ridge or valley in the energy landscape—the system naturally follows the path without being forced sideways.
-The simplest dynamics for evolving a curve toward this MEP is given by:
-Here is a Lagrange multiplier term enforcing the parameterization constraint, typically equal arc-length parameterization where (constant) for all .
-In practice, the string is discretized into images and evolved through a two-step procedure that requires explicit reparameterization steps:
-1. Evolution step:
-2. Reparameterization step to enforce equal spacing:
-This two-step procedure converges to a discretized MEP satisfying each image’s local orthogonality condition .
-2.2.2 Chain-of-States Methods
-An alternative perspective views MEP finding as optimization of a chain-of-states (Pratt, 1986; Elber & Karplus, 1987; Ulitsky & Elber, 1990; Jónsson et al., 1998; Henkelman et al., 2000; Gillilan & Lilien, 2004; Sheppard et al., 2008).
-Given a sequence with and , the objective function is defined as:
-where and is a spring constant.
-The spring term enforces the continuity of the path while allowing flexibility in the discretization.
-However, this formulation introduces a well-known limitation: the spring forces tend to cause "corner cutting," where the path takes shortcuts through high-energy regions rather than following the low-energy valleys representing physically meaningful transitions.
-Although variants of chain-of-states methods exist that avoid this limitation, they come at an increased computational cost.
+The simplest dynamic for evolving toward this MEP is the following:
+where the Lagrange multiplier enforces parameterization constraints.
+However, in practice, implementations alternate between: (1) evolving images along the string using energy gradients , and (2) reparameterization to maintain equal spacing:
+This effectively enforces the same constraint as the Lagrange multiplier (E & Vanden-Eijnden, 2010).
+2.2.3 Computational Limitations
+These classical approaches share fundamental optimization limitations stemming from their discrete optimization in Cartesian molecular coordinates (Ovchinnikov et al., 2011).
+Chain-of-states methods require careful tuning of the spring constant and face corner cutting, while string methods demand complex reparameterization procedures (Lindgren et al., 2019).
+Both struggle with initial transition guesses (Ovchinnikov et al., 2011), local minima trapping, and limited parallelization that restricts applications to smaller systems, as each image is separately parameterized.
+Our neural representation addresses these challenges by encoding continuity architecturally, simplifying optimization to a single energy-based objective while leveraging optimization in neural parameter space to reduce local minima susceptibility.
 3 Experiments
-We evaluate our method on two molecular systems: the AIB9 peptide and Bovine Pancreatic Trypsin Inhibitor (BPTI).
+We evaluate our method on two molecular systems: the AIB9 peptide and the Bovine Pancreatic Trypsin Inhibitor (BPTI).
 These systems represent different scales of complexity and serve to validate our method’s ability to discover physically meaningful transition paths.
-In Appendix A.2, we compare and benchmark our method with the related approaches outlined in (Du et al., 2024; Lee et al., 2025).
-In Appendix B.1, we present ablation studies analyzing the impact of different embedding strategies, interpolation methods, neural network architectures, and regularization approaches.
+In Appendix B, C and D, we show extensive benchmarks of our method with the related approaches outlined in (Du et al., 2024; Lee et al., 2025) and classical MEP approaches, namely chain-of-states, NEB, and the string method, and demonstrate that we outperform all other methods in terms of both mean and maximum path energy.
+In Appendix F, we present ablation studies analyzing the impact of our architectural choices on MEP quality.
 3.1 AIB9 Transition Path Discovery
 We first examine the AIB9 system, a 9-residue artificial protein with 129 atoms that exhibits two well-defined metastable states, making it an ideal test case for transition path methods.
-This system effectively bridges abstract models and complex proteins: it is small enough for extensive reference simulations yet complex enough to exhibit realistic conformational changes; its energy landscape is well-characterized; and its transitions are easily visualized through dihedral angle projections.
+It is small enough for extensive reference simulations yet complex enough to exhibit realistic conformational changes, and its transitions are easily visualized through dihedral angle projections.
+We therefore also use it for the majority of benchmarking and ablation experiments.
 Unlike artificial toy problems, AIB9 allows direct validation with actual biomolecular force fields.
-We use the AMBER ff15ipq-m force field for protein mimetics (Bogetti et al., 2020) implemented in DMFF (Wang et al., 2023), and we generate multiple MEPs between the system’s two metastable states.
+We use the AMBER ff15ipq-m force field for protein-mimetics (Bogetti et al., 2020) implemented in DMFF (Wang et al., 2023), and we generate multiple MEPs between the system’s two metastable states.
 Figure 2 shows the discovered transition paths projected onto the central residue’s - dihedral angle space.
 The method identifies multiple physically plausible pathways between the states, which are also observed with long MD simulations.
-For each path, different conformations of the states were picked as start and end points; this, along with the neural network’s random initialization, leads to the discovery of multiple different transitions.
-Notably, the discovered MEPs align well with the free energy landscape despite our choice of action functional without the trajectory entropy term and the assumption of overdamped dynamics, which may not be ideal for this vacuum system, but is less problematic for hydrated systems.
+For each path, different conformations in the states were picked as start and end points; this, along with the neural network’s random initialization, leads to the discovery of multiple different transitions.
+Notably, the discovered MEPs align well with the free energy landscape despite our choice of an action functional without the trajectory entropy term and the assumption of overdamped dynamics, which may not be ideal for this vacuum system, but is less problematic for hydrated systems.
 4 BPTI Conformational Change Pathway
 Bovine Pancreatic Trypsin Inhibitor (BPTI) is a clinically significant protein used as an anticoagulant in medical procedures.
-BPTI is 58 residues long (892 atoms) and has complex conformational changes involving disulfide bond rearrangements paired with backbone changes.
-To investigate BPTI’s conformational transitions, we use as reference structures five key states from a landmark molecular dynamics (MD) simulation performed by D.E.
+BPTI is 58 residues long (892 atoms) and exhibits complex conformational changes involving disulfide bond rearrangements paired with backbone changes.
+To investigate BPTI’s conformational transitions, we use as reference structures five states from a landmark molecular dynamics (MD) simulation performed by D.E.
 Shaw Research on their Anton supercomputer (Shaw et al., 2010).
-These publicly available snapshots provide rare benchmark structures for validating our approach to BPTI’s transition mechanisms, making it an excellent larger system to test the method.
-Full details of our BPTI system preparation, explicit solvent shell construction, and hybrid neural network architecture for handling water molecules are provided in Appendix B.2.
+These publicly available snapshots provide rare benchmark structures for validating our approach to BPTI’s transition mechanisms, making it an excellent system to test the method on a larger scale.
+Full details of our BPTI system preparation and hydration are provided in Appendix G.
 We generate multiple transition MEPs connecting two BPTI states from the D.E.
@@
 Remarkably, as shown in Figure 3, our optimized path passes through all intermediate snapshots, indicating that it successfully discovers the same transition mechanism as the computationally intensive MD simulation.
-We matched their simulation conditions by incorporating explicit water molecules in our model to ensure a meaningful comparison with the Anton simulation.
-We generated water shells extending 14Å from any protein atom for both the initial and final BPTI conformations.
-This increases the number of atoms from 892 to 8,300.
+We matched their simulation conditions by incorporating explicit water molecules to ensure a meaningful comparison with the Anton simulation.
+We hydrate both the initial and final BPTI conformations, as done in a conventional MD system setup.
+This increases the number of atoms from 892 to 3,500.
 To handle water molecules during transitions, we pair proximal water molecules between start and end states using the Hungarian algorithm (Kuhn, 1955).
-We then model their movement using simple parameterized splines, rather than directly incorporating all water coordinates into the neural network.
-This hybrid neural-spline approach allows us to include explicit solvent effects while keeping computational costs low.
-The specific details of our water shell construction and matching algorithm are described in Appendix B.2.
 The system was modeled using the AMBER99sb-ILDN force field (Lindorff-Larsen et al., 2010) with the TIP3P water model (Jorgensen et al., 1983; Neria et al., 1996).
@@
 This represents a six orders of magnitude reduction in computational effort.
-While our approach does not provide the same statistical ensemble information as equilibrium MD, it efficiently captures the key transition pathway, which is often the primary mechanistic insight sought by MD practitioners.
-For most scientifically interesting systems—typically large biomolecular complexes—equilibrium sampling of the complete conformational space or transition path ensemble remains computationally prohibitive, making identification of dominant pathways particularly valuable.
-It’s worth noting that even the reference millisecond-scale simulation may not have fully achieved equilibrium sampling of the conformational space.
-Appendix B.2.7 provides a detailed computational performance comparison.
+While our approach does not provide the same statistical ensemble information as the reference MD, it efficiently captures the key transition pathway.
+It is worth noting that even the reference millisecond-scale simulation has not achieved equilibrium sampling, as it is comprised of a single transition.
+Appendix I provides a detailed computational performance comparison.
+5 Comparative Evaluation and Ablation Studies
+To systematically validate our approach, we conducted extensive benchmarks against established methods and ablation studies examining our architectural choices across three molecular systems of increasing complexity.
+We evaluate path quality using two key metrics: the maximum energy along the path (indicating transition state barriers), the arc-length adjusted mean energy (accounting for how quickly paths traverse high-energy regions), and the standard deviation of this quantity for non-deterministic methods.
+The full results for the benchmarks are presented in sections B, C, and D, and the ablations are discussed in section F of the appendix, respectively.
+Benchmark Results: AdaPath consistently outperforms classical approaches (Chain-of-States, String Method, NEB) and deep learning approaches like Doob’s Lagrangian in both computational efficiency and path quality.
+For the alanine dipeptide, we achieve superior energy barriers (-86.6 vs.
+-71.1 kJ/mol for Chain-of-States) while requiring fewer force evaluations (6.4e4 vs.
+2.4e5).
+Here and on other systems, Chain-of-States emerged as the best-performing classical method.
+We note that Doob’s Lagrangian targets the more ambitious objective of modeling the entire transition path ensemble rather than discovering a single MEP, making direct performance comparisons inherently limited.
+This performance advantage scales to larger systems: on AIB9 (129 atoms), AdaPath reaches -644 kJ/mol maximum energy compared to Chain-of-States’ -461 kJ/mol and substantially outperforms all other tested methods (Appendix C).
+On the BPTI system with explicit solvent (3,500 atoms), classical methods either fail to converge and produce erratic paths that bypass intermediate reference structures from the landmark D.E.
+Shaw simulation, as shown qualitatively in Appendix D.
+Architectural Validation: Ablation studies confirm our core design choices (Appendix F).
+Parameter sharing proves essential for scalability; standard MLPs fail catastrophically on large systems, with even billion-parameter variants showing poor performance on BPTI.
+Our optimizer analysis reveals that Muon maintains stable path continuity, while Adam exhibits erratic ”teleportation” behavior that rapidly jumps between states without traversing transition regions in small systems.
+The logarithmic energy transformation proves critical for training stability, reducing energy variance by orders of magnitude.
 6 Limitations
-While our feedforward neural architecture can model conformational transitions, it could benefit from incorporating domain-specific structural priors to improve training efficiency and convergence speed.
-Furthermore, physical force fields introduce challenges through high-energy spikes that can cause training instability and divergence in some runs.
-Our gradient clipping approach, outlined in B.2.5, largely mitigates this issue but doesn’t universally prevent training failures.
-The method has been validated primarily on small to medium-sized protein systems, and scaling to larger multi-domain proteins could require additional optimization strategies.
-Our approach is also constrained to systems with suitable differentiable force fields, but this limitation diminishes as more implementations become available.
-Finally, while computationally efficient, our method focuses on identifying MEPs rather than sampling the full statistical ensemble of transition pathways, representing a trade-off between efficiency and exhaustively sampling a system.
-However, as noted in the BPTI section 4, this mechanistic insight is often a key objective for MD practitioners studying complex biomolecular transitions.
+While our approach demonstrates strong performance on protein systems up to 3,500 atoms, several limitations should be noted.
+First, our method is constrained to systems with suitable differentiable force fields, though this limitation is diminishing as more implementations become available.
+Second, our method produces single representative MEPs rather than sampling the complete statistical ensemble of transition pathways.
+While this represents a trade-off between computational efficiency and exhaustive sampling, the mechanistic insights provided are often the primary objective for practitioners studying biomolecular transitions.
 7 Conclusion
-We demonstrated that reformulating MEP discovery as a continuous neural optimization problem enables efficient discovery of molecular transition mechanisms with unprecedented computational efficiency.
-Our approach transforms the complex Onsager-Machlup action functional into a straightforward energy-based loss function that eliminates the need for force and path gradient computations required by related methods, which are known to induce optimization challenges in PINNs.
-This work provides a practical tool for molecular dynamics practitioners that captures essential mechanistic insights without requiring specialized hardware or extensive simulation time.
-While our current implementation shows promising results, several concrete directions for improvement remain.
-Future work should explore neural architectures that incorporate domain-specific structural priors beyond simple MLPs with sinusoidal/Fourier embeddings.
-Related work by Ramakrishnan et al.
-(2025) investigates non-uniform sampling schemes along the path that might further aid convergence for paths that constitute dramatic rearrangements.
-Integrating pretrained generative models suggested by Raja et al.
-(2025) could provide better initial transition guesses, potentially accelerating convergence by starting optimization from conformations that already respect molecular geometry.
-Extending the method to handle larger multi-domain proteins is also a frontier of great scientific interest, which might bring new challenges.
-Future work could also explore ways to estimate free energies and reaction rates from our discovered paths.
-7 Broader Impacts
-Our research democratizes access to the study of larger molecular systems by enabling transition path studies on standard hardware.
-This may accelerate drug discovery and protein engineering while contributing to sustainable computing through drastically reduced energy consumption.
-We see no direct negative societal impacts, though we acknowledge that accelerated protein engineering capabilities could be misused in the future, and we encourage responsible use with appropriate experimental validation.
-Appendix A Technical Appendices and Supplementary Material
+We demonstrated that reformulating MEP discovery as a continuous neural optimization problem, combined with a scalable neural architecture, enables efficient discovery of molecular transition mechanisms for large biomolecular systems in explicitly modeled solvent.
+Our approach makes two key contributions: deriving a simple energy-based loss function directly from the Onsager-Machlup action functional, and introducing AdaPath.
+This architecture scales favorably with system size via parameter sharing across the atom dimension.
+By encoding path continuity through architectural design rather than explicit loss terms, we sidestep known optimization challenges in PINNs and multi-loss training.
+Several promising directions for future development emerge from this work.
+The scaling properties of AdaPath suggest the method could be extended to even larger biomolecular systems, such as membrane proteins or multi-domain complexes that routinely contain tens of thousands of atoms in MD simulations, not counting water.
+Our discovered MEPs could serve as guidance paths for enhanced sampling methods, enabling free energy calculations through escorted Jarzynski-like estimators (Jarzynski, 1997; Vaikuntanathan & Jarzynski, 2011) by using the MEP as a basis for a non-equilibrium pulling protocol between stable states.
+The architecture and loss function may also prove valuable for other non-biological physical systems where rare events might be of interest.
+Appendix A Appendix
 A.1 Extended Derivation of the energy-based Loss Function
@@
 We also make several important approximations during the derivation:
-We use a first-order Taylor expansion to relate energy differences to gradient magnitudes
+We use a first-order Taylor expansion to relate energy differences to gradient magnitudes, which requires displacement vectors to be sufficiently small
 We introduce stochastic sampling to estimate the continuous energy integral
-For a MEP, the displacement tends to align either parallel or anti-parallel with the gradient direction (i.e., for path segments)
+For a MEP, the displacement tends to align either parallel or antiparallel with the gradient direction (i.e., for path segments)
 We use the triangle inequality to establish an upper bound on the sum of absolute energy differences
 While the last two approximations do not change the location of the global minimum, they do affect the optimization landscape before that minimum is reached.
+Critically, when the Taylor approximation breaks down due to large displacement vectors, the resulting energy-based loss can admit pathological solutions where paths remain in stable states and then rapidly ”teleport” to avoid high-energy transition regions.
 A.1.1 From the Onsager-Machlup Action to a geometric Line Integral
@@
 By the Cauchy-Schwarz inequality:
-with equality when is parallel or anti-parallel to , with the direction determined by whether the path is ascending or descending the energy landscape.
+With equality when is parallel or antiparallel to , with the direction determined by whether the path is ascending or descending the energy landscape.
 This is precisely the defining characteristic of the MEP - a path that follows the potential energy landscape’s gradient while avoiding high-energy regions.
@@
 Using Taylor’s theorem, we can relate energy differences to gradient magnitudes:
-where is the remainder of the order and is the angle between the gradient and displacement vectors.
+where is the angle between the gradient and displacement vectors.
+Critically, this Taylor expansion approximation is only valid when is sufficiently small for the higher-order term to be negligible.
+When displacement vectors become large, the approximation breaks down.
+It can lead to pathological solutions, where the optimization finds paths that avoid high-energy transition regions by rapidly ”teleporting” between stable states, rather than discovering smooth, minimum-energy routes.
 For a path following the minimum-energy trajectory, the gradient aligns with or against the path direction ( as , depending on whether the path is ascending or descending the energy landscape), giving at the minimum of the optimization:
 This yields an approximation of the geometric action:
-This forms a surrogate objective with the same minimum as our original action.
+This forms a surrogate objective with the same minimum as our original action, provided the Taylor approximation remains valid throughout optimization.
 A.1.3 Establishing the energy-based Loss Function
@@
 The resulting path follows the potential energy landscape’s gradient while avoiding high-energy regions, representing the most probable transition mechanism in the overdamped limit.
-A.2 Connection between Doob’s Lagrangian and MEP Optimization
-We elaborate on the theoretical connections between our energy-based optimization approach and the framework of Doob’s h-transform for sampling transition paths (Du et al., 2024).
-This comparison is particularly relevant as both methods employ neural networks to represent molecular state transitions—either as a minimum-energy path in our case or the full transition path ensemble in theirs.
-Notably, the parametrization of the mean trajectory in Du et al.
-resembles our MEP representation, differing primarily in our use of blending functions and spatial embeddings for the neural network component.
-Doob’s Lagrangian formulation considers the following action functional:
-This functional quantifies the cost of controlling a stochastic process to achieve desired endpoint conditions, subject to the system’s energy landscape.
-Intuitively, it measures the amount of "effort"/action needed to steer the dynamics from starting state A to target state B, with smaller values indicating more probable transition paths.
-In this formulation, is the probability density at time given boundary conditions, parameterized as a Gaussian , and is the control vector field satisfying , with representing the reference drift for overdamped dynamics.
-This action functional describes a Wasserstein Lagrangian flow as formalized in (Neklyudov et al., 2023) and (Neklyudov et al., 2024).
-In this framework, one can learn stochastic dynamics from samples, thereby learning a process that defines a time-dependent density evolving from initial to final states.
-The key difference with our PINN-MEP approach is that we minimize the action of a single deterministic path rather than a full distribution of paths.
-Mathematically, this corresponds to working with a single trajectory through configuration space rather than an ensemble of possible trajectories.
-This simplification is reflected in our loss function, which contains one fewer expectation than the full Doob’s Lagrangian.
-A.2.1 Comparison with Doob’s Lagrangian Method
-We provide a direct comparison between our PINN-MEP approach and the method of Du et al.
-in terms of computational efficiency, as evaluated by the number of force field/energy function calls, and the maximum energies observed in the paths, corresponding to the peak of the potential barrier.
+However, as noted above, the validity of this approach depends critically on maintaining sufficiently small displacement vectors during optimization to ensure the underlying Taylor approximation remains valid.
+Appendix B Benchmarks: Alanine Dipeptide
+We provide a direct comparison between our AdaPath method, Du et al.
+(2024), Lee et al.
+(2025), as well as classical, well-established methods like the string method, chain-of-states, and nudged elastic band, in terms of computational efficiency, as evaluated by the number of force field/energy function calls, and the maximum energies observed along the paths, corresponding to the peak of the potential barrier, both as a mean across five seeds as well as the lowest value achieved across all runs.
+We further compute the arc-length-adjusted mean energy along the path.
+We chose this metric as naive averaging along a discretized path does not account for how fast a path is traversing a high-energy region; it could, therefore, shortcut a region of high energy but have a very low non-adjusted energy.
 To provide a rigorous comparison, we conducted benchmarks using experimental conditions identical to those reported by Du et al.
-for alanine dipeptide: AMBER14 force field, 300K temperature.
+for alanine dipeptide: AMBER14 force field, 300 K temperature, and also referred to their results when applicable, but adjusted the significant figure notation.
 Table 1 presents these results alongside those reported in Du et al.
@@
 (2025).
-The only exception is that we do not evaluate trajectory probability, as the assumption of optimization over a non-physical path length makes the calculation of this quantity impossible.
-Method | States | # Evaluations | Max Energy | MinMax Energy |
-MCMC variable length | CV | 21.02M | 740.70 695.79 | 52.37 |
-MCMC* | CV | 1.29B | 288.46 128.31 | 60.52 |
-MCMC variable length | relaxed | 187.54M | 412.65 334.70 | 26.97 |
-MCMC | relaxed | 10B | N/A | N/A |
-MCMC variable length | exact | 10B | N/A | N/A |
-MCMC | exact | 10B | N/A | N/A |
-Doob’s Cartesian | exact | 38.40M | 726.40 0.07 | 726.18 |
-Doob’s Cartesian, 2 Mixtures | exact | 51.20M | 709.38 162.37 | 513.72 |
-Doob’s Cartesian, 5 Mixtures | exact | 51.20M | 541.26 278.20 | 247.96 |
-Doob’s Internal | exact | 38.40M | -14.62 0.02 | -14.67 |
-Doob’s Internal, 2 Mixtures | exact | 51.20M | -15.38 0.14 | -15.54 |
-Doob’s Internal, 5 Mixtures | exact | 51.20M | -15.50 0.31 | -15.95 |
-Doob’s Seq2Seq | exact | - | 1505.6 0.45 | 245.05 0.02 |
-w/ Temperature Annealing | exact | - | 1583.18 0.3 | 592.13 0.26 |
-w/ Interatomic Interpolation | exact | - | 1601.57 0.53 | 3.46 0.03 |
-MaxLL | exact | - | 1532.45 | 615 |
-w/ Temperature Annealing | exact | - | 1599.03 | 233 |
-w/ Interatomic Interpolation | exact | - | 1545.32 | 619 |
-PINN-MEP | exact | 15700 | -67.22 3.55 | -70.78 |
-PINN-MEP requires approximately 3261× fewer energy function evaluations (15.7K vs.
-51.20M) compared to the Doob’s Lagrangian approach while achieving a 4.44 times lower maximum energy barrier.
-The lower number of required evaluations is likely due to the computation of the loss, which requires sampling their path over both , as we do as well, but also the width of the path.
-Thus, the tradeoff of sacrificing full ensemble characterisation is expected to come with computational efficiency.
-Appendix B Comparison with other related Methods
+The only exception is that we do not evaluate trajectory probability, as the assumption of optimization over a non-physical path length makes calculating this quantity nonsensical.
+Furthermore, the arc-length-adjusted energy quantities are not applicable to Doob’s method, as they do not model a single path but rather the entire transition path ensemble.
+For the classical methods, we conducted extensive hyperparameter sweeps to ensure fair comparison and optimal performance.
+Chain-of-States was optimized across path discretizations (20-60 images), spring constants (1e-7 to 1e-1), and learning rates (5e-2 to 1e-1), with the best configuration using 40 images, a spring constant of 1e-6, and the Adam optimizer with a learning rate of 5e-2.
+The String Method was evaluated across image counts (20-60), timesteps (0.1-4000), and reparameterization frequencies (1-10 steps), achieving optimal performance with 40 images, timestep 3900, and reparameterization every three steps.
+NEB was tested across various path discretizations (20-60 images), spring constants (1e-7 to 1e-1), and learning rates (1e-6 to 1e-1), with the best results obtained using 60 images, a spring constant of 2e-5, and the Adam optimizer with a learning rate of 1e-5.
+These configurations represent the optimal settings identified through systematic exploration of the hyperparameter space for each method.
+All classical approaches we study are deterministic, and hence are only run once, unlike the stochastic approaches, which are run five times.
+Since the field lacks standardized benchmarks, we believe that building on the work from (Du et al., 2024) could be a helpful step toward having a more standardized approach to methods aiming to solve for MEPs, the transition path ensemble, or other similar approaches.
+We want to reiterate that (Du et al., 2024) is more ambitious in its sampling objectives, and therefore, one-to-one comparisons are not entirely appropriate.
+However, the transition state energy, as represented by the Max Energy entry, should be consistent between MEP approaches and methods that aim to sample the entire transition ensemble.
+| Method | # Evals | Max Energy | MinMax Energy | Mean Energy |
+| MCMC variable length | 2.1e7 | N/A | ||
+| MCMC* | 1.3e9 | N/A | ||
+| MCMC variable length | 1.9e8 | N/A | ||
+| MCMC | 1e10 | N/A | N/A | N/A |
+| MCMC variable length | 1e10 | N/A | N/A | N/A |
+| MCMC | 1e10 | N/A | N/A | N/A |
+| Doob’s Cartesian | 3.8e7 | N/A | ||
+| w/ 2 Mixtures | 5.1e7 | N/A | ||
+| w/ 5 Mixtures | 5.1e7 | N/A | ||
+| Doob’s Internal | 3.8e7 | N/A | ||
+| w/ 2 Mixtures | 5.1e7 | N/A | ||
+| w/ 5 Mixtures | 5.1e7 | N/A | ||
+| Doob’s Seq2Seq | – | N/A | ||
+| w/ Annealing | – | N/A | ||
+| w/ Interpolation | – | N/A | ||
+| MaxLL | – | N/A | ||
+| w/ Annealing | – | N/A | ||
+| w/ Interpolation | – | N/A | ||
+| Chain-of-States | 2.4e5 | |||
+| String Method | 3.3e6 | |||
+| NEB | 7.7e6 | |||
+| AdaPath | 6.4e4 |
+AdaPath requires approximately fewer energy function evaluations (6.4e4 vs.
+2.4e5) compared to the best classical method (Chain-of-States) while achieving superior energy barriers ( vs.
+kJ/mol for the minimum maximum energy).
+Compared to the Doob’s Lagrangian approach, AdaPath requires approximately fewer energy function evaluations (6.4e4 vs.
+5.1e7) while achieving significantly lower maximum energy barriers.
+The classical methods demonstrate substantial variation in computational efficiency, with Chain-of-States providing the best computational performance among baseline approaches in terms of function evaluations.
+However, AdaPath still requires fewer evaluations while achieving better energy minimization.
+The lower number of required evaluations for AdaPath is likely due to the efficient neural representation compared to discrete path sampling approaches used by classical methods, and the large number of images that these methods need to function well.
+Appendix C Benchmarks: AIB9
+We conducted similar benchmarks on the AIB9 peptide system, a 129-atom artificial protein that exhibits well-defined metastable states and realistic conformational changes.
+This system serves as an intermediate complexity benchmark between simple dipeptides and larger protein systems.
+For the classical methods, we performed comprehensive hyperparameter optimization: Chain-of-States was evaluated across path discretizations (20 to 60 images), spring constants (1e-6 to 1e-1), and learning rates (1e-3 to 1e-1), achieving optimal performance with 40 images, spring constant 1e-5, and Adam optimizer with learning rate 1e-2.
+String Method was tested across image counts (20 to 60), timesteps (1 to 20), and reparameterization frequencies (1 to 15 steps), with best results using 60 images, timestep 10, and reparameterization every three steps.
+NEB was optimized across path discretizations (20 to 60 images), spring constants (1e-7 to 1e-1), and learning rates (1e-6 to 1e-1), yielding an optimal configuration with 60 images, spring constant 1e-3, and Adam optimizer with learning rate 1e-5.
+For the Doob’s Lagrangian method, we adopted the settings reported for the Chignolin system in Du et al.
+(2024), as it represents a similar-sized protein system.
+Due to observed training instability, we added gradient clipping (1e-2) to stabilize optimization.
+We also found the original parameterization to be slightly under-parameterized for this system and increased the MLP width to 768 per layer across five layers.
+While the resulting energies appear high, they are consistent with the 3000 kJ/mol values reported for Chignolin in the original work, suggesting this energy scale is characteristic of the method’s behavior on protein systems of this size.
+| Method | # Evals | Max Energy | MinMax Energy | Mean Energy |
+| Chain-of-States | 2.9e6 | |||
+| String Method | 3.1e6 | 1.3e4 | 1.3e4 | 3.2e3 |
+| NEB | 4.1e6 | 2.7e3 | 2.7e3 | |
+| Doob’s Lagrangian | 1.2e7 | 5.8e3 | N/A | |
+| AdaPath | 4.7e6 |
+The results demonstrate AdaPath’s superior performance compared to both classical methods and the Doob’s Lagrangian approach on this complex system.
+AdaPath achieved the best overall energy minimization, with a minimum maximum energy of kJ/mol, substantially outperforming Chain-of-States ( kJ/mol), String Method (1.3e4 kJ/mol), NEB (2.7e3 kJ/mol), and Doob’s Lagrangian (5.8e3 kJ/mol).
+The arc-length-adjusted mean energies further highlight AdaPath’s advantages, achieving kJ/mol.
+Notably, AdaPath demonstrated both superior energy minimization and significantly improved consistency across multiple seeds, with 3 out of 5 seeds achieving excellent results below kJ/mol, while Doob’s Lagrangian showed high variability with a standard deviation of 7.8e3 kJ/mol.
+The substantial energy differences between methods highlight the importance of the optimization objective and neural architecture design for effective transition path discovery.
+Appendix D Benchmarks: BPTI
+We conducted benchmarks on the BPTI (Bovine Pancreatic Trypsin Inhibitor) system to evaluate our AdaPath method against classical transition path discovery approaches on a significantly larger and more complex biomolecular system, BPTI.
+The system setup, including explicit solvation, ionic strength, and force field parameters, is detailed in Appendix G.
+For the classical methods (Chain-of-States, String Method, and NEB), we employed the same hyperparameter configurations that proved optimal for the AIB9 system rather than conducting comprehensive sweeps, as the computational cost of extensive hyperparameter optimization on a system of this size would be prohibitive.
+Specifically, we used the following methods: Chain-of-States with 40 images, a spring constant of 1e-5, and the Adam optimizer with a learning rate of 1e-2; String Method with 60 images, a timestep of 10, and reparameterization every three steps; and NEB with 60 images, a spring constant of 1e-3, and the Adam optimizer with a learning rate of 1e-5.
+For AdaPath, we used the training parameters specified in Appendix G, which is a more parameter-efficient variant of the model that performed best on the smaller AIB9 system.
+| Method | # Evals | Max Energy | MinMax Energy | Mean Energy |
+| Chain-of-States | 1.4e6 | |||
+| String Method | 3.8e6 | 1.4e4 | 1.4e4 | 4.3e3 |
+| NEB | 1.1e6 | 1.6e5 | 1.6e5 | 2.3e5 |
+| AdaPath | 1.9e5 |
+The results demonstrate AdaPath’s superior performance on this complex, explicitly solvated system.
+Figure 4 shows that among classical approaches, Chain-of-States produces the most reasonable pathway but fails to navigate through all intermediate reference structures from the D.E.
+Shaw simulation.
+The String Method fails to converge to a good solution, while NEB displays a jagged path, a known issue for this method in high-dimensional systems (Zhang et al., 2016).
+In contrast, AdaPath successfully discovers transition paths that pass through all intermediate reference structures, as demonstrated in Figure 3, while achieving significantly lower maximum energy barriers.
+Appendix E Other related Methods
 While our primary benchmark comparison focuses on the Doob’s Lagrangian method due to similarity in task and system setup, examining how our approach relates to other methods for transition path discovery/MEP generation is also informative.
-Here we present results from two related approaches in the literature, though we emphasize that direct numerical comparisons should be treated with caution due to differences in force fields, optimization objectives, and units used.
+Here, we present results from two related approaches in the literature.
+However, we emphasize that direct numerical comparisons should be treated with caution due to differences in force fields, optimization objectives, and units used.
 | Method | # Evaluations | Max Energy (eV) |
-NEB | 8063 | 11.0014 |
-INR | 3046 | 0.4070 |
+| NEB | 8.1e3 | |
+| INR | 3.0e3 |
 | Method | CVs | # Evaluations | Runtime / Path |
-MCMC (Two-Way Shooting) | No | 1B | 100 hours |
-Metadynamics | Yes | 1M | 10 hours |
-OM Opt. (Diffusion Model) | No | 10K | 50 min |
-Table 2 shows results from Ramakrishnan et al.
-(2025) comparing Nudged Elastic Band (NEB) with their Implicit Neural Representation (INR) approach for MEP discovery.
-Both our PINN-MEP approach and their INR method utilize lightweight, preexisting molecular dynamics force fields, though we employ different such force fields.
-Their approach requires a more complex loss function that evaluates both energy and forces along with path parameterization and its gradient (path velocity), but needs fewer force field evaluations than our method.
-It is important to note that the "Max Energy" reported in Table 2 is measured in electron volts (eV) and appears to be adjusted for ground state energy reference, which makes direct energy comparisons with our method challenging.
+| MCMC (Two-Way Shooting) | No | 1e9 | 100 hours |
+| Metadynamics | Yes | 1e6 | 10 hours |
+| OM Opt. (Diffusion Model) | No | 1e4 | 50 min |
+Table 4 presents results from Ramakrishnan et al.
+(2025) comparing the Nudged Elastic Band (NEB) method with their Implicit Neural Representation (INR) approach for MEP discovery.
+Both our AdaPath approach and their INR method utilize lightweight, preexisting molecular dynamics force fields, though we employ different such force fields.
+Their approach requires a more complex loss function that evaluates both energy and forces along with path parameterization and its gradient (path velocity).
+However, it needs fewer force field evaluations than our method.
+It is important to note that the ”Max Energy” reported in Table 4 is measured in electron volts (eV) and appears to be adjusted for ground state energy reference, which makes direct energy comparisons with our method challenging.
 The difference in force fields used further complicates any direct energy value comparisons between methods.
@@
 Their approach requires evaluating both the pretrained model (as a surrogate force field) and its gradient for their loss function, which is more computationally intense due to the high number of parameters of the pretrained model as compared to a physics-derived force field.
-These tables corroborate our findings in the Doob’s Lagrangian comparison: methods optimizing for a single MEP consistently require orders of magnitude fewer force field evaluations than approaches that characterize the full transition path ensemble.
-B.1 Ablation Studies
-This section presents a series of ablation studies examining the impact of various components of our PINN-MEP approach.
+These tables corroborate our findings in the Doob’s Lagrangian comparison: methods that optimize for a single MEP consistently require orders of magnitude fewer force field evaluations than approaches that characterize the complete transition path ensemble.
+E.1 Connection between Doob’s Lagrangian and MEP Optimization
+We elaborate on the method of Doob’s h-transform for sampling transition paths (Du et al., 2024).
+This explanation is relevant as both methods employ neural networks to represent molecular state transitions, either as a minimum-energy path in our case or the complete transition path ensemble in theirs.
+Notably, the parametrization of the mean trajectory in Du et al.
+resembles our MEP representation, differing primarily in our use of blending functions and spatial embeddings for the neural network component.
+Doob’s Lagrangian formulation considers the following action functional:
+This functional quantifies the cost of controlling a stochastic process to achieve desired endpoint conditions, subject to the system’s energy landscape.
+Intuitively, it measures the amount of ”effort”/action needed to steer the dynamics from starting state A to target state B, with smaller values indicating more probable transition paths.
+In this formulation, is the probability density at time given boundary conditions, parameterized as a Gaussian , and is the control vector field satisfying , with representing the reference drift for overdamped dynamics.
+This action functional describes a Wasserstein Lagrangian flow as formalized in (Neklyudov et al., 2023) and (Neklyudov et al., 2024).
+In this framework, one can learn stochastic dynamics from samples, thereby learning a process that defines a time-dependent density evolving from an initial to a final state.
+The difference with our approach is that we minimize the action of a single deterministic path rather than a complete distribution of paths.
+This simplification is reflected in our loss function, which contains one fewer expectation than the full Doob’s Lagrangian.
+Appendix F Ablation Studies
+This section presents a series of ablation studies examining the impact of various components of our AdaPath approach.
 These experiments help validate our design choices and provide insights into the importance of different architectural and training elements.
 We conduct these ablation studies on the AIB9 system because Alanine dipeptide, as a benchmark system, is too simple to provide meaningful insights into the performance of our method.
-For each ablation study, we report the lowest maximum energy of the path among all training runs, the mean and standard deviation of the energy, and the mean number of iterations to reach the lowest maximum energy.
-We run each setting in the ablations 15 times for a maximum of 25,000 iterations.
-We use a batch size of 1, where each batch contains 32 values of the progress coordinate for most cases.
-The exception is when we evaluate the chain-of-states-like regularization, which minimizes the difference between network outputs for adjacent values of ; in this case, the effective batch size is 32, since there are no interactions between different network outputs otherwise.
-Due to the nature of physical force fields, slight deviations in the path can lead to large changes in energy, leading to similar settings having very different energies.
-This can limit the ability to compare the performance of different settings.
-However, we find that the results are still directionally informative.
-B.1.1 General Settings for Ablation Studies
+For each ablation study, we report the lowest maximum energy of the path among all training runs, the mean and standard deviation of the arc-length adjusted energy, and the mean number of iterations to reach the lowest maximum energy.
+We run each setting in the ablations 5 times.
+We use a batch size of 1, where each batch contains 16 values of the progress coordinate for most cases.
+F.1 General Settings for Ablation Studies
 All ablation studies are conducted using the following base configuration, with each study varying only specific parameters while keeping all others fixed:
@@
 Training Configuration:
-Number of iterations: 25,000 per run
-Number of independent runs: 15 per parameter setting
-Evaluation frequency: Every 100 training steps
-Optimizer: Adam with optax default parameters
-Learning rate:
+Number of iterations: 100,000 per run
+Number of independent runs: 5 per parameter setting
+Evaluation frequency: Every 2000 training steps
+Optimizer: Muon as implemented in Optax.
+Learning rate: 1e-2
 Path Sampling Parameters:
 Batch size: 1
-Frames per batch: 32 points sampled along the progress coordinate
-Spring constant: (no chain-of-states regularization)
-Interpolation method: Cubic Hermite for blending functions
+Frames per batch: 16 points sampled along the progress coordinate
+Interpolation method: Linear
 Network Architecture (Base Configuration):
-Network depth: 6 hidden layers
-Width factor: 4 (network width scales as number of atoms)
-Embedding dimension: Automatically set to
+Network depth: 6 MLP blocks.
+MLP Expansion Factor: 4, meaning that each MLP block, consisting of two linear layers, has four times the hidden dimensionality as input and output dimensionality.
+Atom Embedding Dimension: 32.
 Time embedding: Sinusoidal with maximum period of
-Activation function: ReLU throughout the network
-Weight initialization: Glorot uniform for hidden layers, zero initialization for final layer
-Bias initialization: Default for hidden layers, zero initialization for final layer
-B.1.2 Embedding Strategy
-We compared sinusoidal time embeddings with different maximum periods against simple linear projections of the progress coordinate.
-Table 4 shows that linear embedding (None) performed best, achieving the lowest mean energy, though sinusoidal embeddings with period also achieved competitive results.
-We find these results at odds with the need for sinusoidal embeddings in the larger BPTI system.
-Perhaps the AIB9 system, which is characterized by one large general conformational rearrangement primarily driven by main-chain dihedral angles, does not need the high frequency features that sinusoidal embeddings provide, but which might be essential for other systems, due to the well-known bias of MLPs toward modeling low frequencies in data.
-Embedding | Max Period | MeanMax Energy () | MinMax Energy | Mean Iterations |
-Linear (None) | – | 591.35 963.72 | -337.36 | 200873 |
-Sinusoidal | 1326.13 2339.35 | -337.36 | 210907 | |
-Sinusoidal | 1528.18 2429.05 | -129.90 | 226290 | |
-Sinusoidal | 1888.66 3604.47 | -337.36 | 187493 | |
-Sinusoidal | 2453.53 3507.47 | -337.36 | 217053 |
-B.1.3 Spring Regularization
-We tested the effect of chain-of-states-like spring constant regularization, as shown in Table 5.
-Interestingly, we found that small amounts of spring regularization ( or ) provided the best results, with achieving the lowest mean energy overall.
-This suggests that a minimal amount of explicit path continuity enforcement can be beneficial when combined with our neural representation, though this is not a strict requirement for the method to work.
-However, adding a spring constant reintroduces intra-batch dependencies, which means that batch sizes should be of a minimum size for this approach to work effectively.
-With low batch sizes, this could lead to the corner cutting effects commonly observed in traditional chain-of-states methods, which might make scaling the addition of spring regularization difficult for larger systems.
-Since we are able to reach the same minimum energy without spring regularization, we recommend against its use, especially when scaling to larger systems where lower batch sizes are preferred.
-Spring Constant | MeanMax Energy () | MinMax Energy | Mean Iterations |
--178.22 260.09 | -337.36 | 153533 | |
--155.41 286.63 | -337.36 | 101600 | |
-803.61 586.95 | 205.74 | 159033 | |
-0.0 | 1460.41 2486.05 | -337.36 | 209133 |
-10528.4 7394.89 | 5065.46 | 155400 | |
-48639.7 12111.7 | 33143.6 | 107267 |
-B.1.4 Network Architecture
-We investigated the effect of network depth and width on transition path quality.
-Table 6 shows that an 8-layer network with a width factor of 4 achieved the best mean energy performance, followed closely by a 6-layer network with width factor 6.
-Depth | Width Factor | MeanMax Energy () | MinMax Energy | Mean Iterations |
-8 | 4 | 25.02 467.23 | -337.36 | 159333 |
-6 | 6 | 72.42 458.72 | -337.36 | 210267 |
-8 | 6 | 165.91 1021.79 | -337.36 | 219433 |
-8 | 2 | 1434.97 3206.06 | -337.36 | 178167 |
-6 | 4 | 1460.34 2486.07 | -337.36 | 209067 |
-6 | 2 | 3621.8 3198.25 | 217.83 | 231200 |
-4 | 6 | 4242.74 3568.51 | 140.81 | 231967 |
-4 | 4 | 5130.75 5758.01 | -328.49 | 228133 |
-4 | 2 | 16200.5 8032.18 | 159.21 | 210367 |
-B.1.5 Interpolation Type
-We examined different interpolation methods for the blending functions that combine the neural network output with the endpoint constraints.
-Table 7 compares linear, cosine, quintic, and cubic Hermite interpolation.
-Cosine interpolation performed best in terms of mean energy, though all methods were able to achieve the minimum energy of -337.36 in at least one run, except for quintic.
-The cubic Hermite interpolation showed higher variance in results.
-The specific formulas for each interpolation type are as follows:
-Linear interpolation:
-Cubic Hermite interpolation:
-Quintic interpolation:
-Cosine interpolation:
-where and are the basis functions that ensure endpoint constraints and , and is the blending function that controls the neural network’s contribution along the path.
-Interpolation Type | MeanMax Energy () | MinMax Energy | Mean Iterations |
-Cosine | 442.51 769.51 | -337.36 | 198720 |
-Linear | 602.61 1057.13 | -337.36 | 219147 |
-Quintic | 967.16 1288.57 | -281.14 | 210227 |
-Cubic Hermite | 1325.49 2339.75 | -337.36 | 212807 |
-B.1.6 AIB9 Network Architecture
-Below are the main details of the AIB9 MEP training runs, as shown in the main paper.
-Model type: Transition MLP with sinusoidal time embedding
-Network depth: 6 hidden layers
-Network width: 4 width factor (scaled relative to the number of atoms)
-Time embedding: Sinusoidal position encoding with maximum period of 1,000,000
-Activation function: ReLU
-B.1.7 Training Parameters
-Optimizer: Adam
-Learning rate:
-Batch size: 4
-Frames per batch: 32
-Spring constant (): 0.0
-Interpolation method: Linear
-All other parameters are identical to the base settings listed in the ablations section B.1.1.
-B.2 BPTI Modeling and Computational Details
-B.2.1 System Preparation
+Activation function: GeLu throughout the network
+Conditioning MLP settings: 3 Layers, with a hidden dimension of 1024 and GeLu activations.
+F.2 Architecture Comparison: AdaPath vs Standard MLP
+To validate our core claim that parameter sharing across atoms provides superior scalability, we compare AdaPath against a standard MLP baseline that directly maps progress coordinate to the complete molecular configuration without parameter sharing.
+This comparison is conducted across all three molecular systems (Alanine dipeptide, AIB9, and BPTI) to demonstrate scaling behavior.
+The standard MLP baseline uses a conventional feedforward architecture with six hidden layers of width determined by 4 times the number of atoms in the system.
+This was found to be suitable for the two small systems.
+Intuitively, this aligns well with deep learning’s collective wisdom of having a higher hidden dimension than the input and output dimensions of the network.
+We only lower this to one time the atom count for the BPTI system, as this already leads to an MLP with 1= billion parameters, as more was not runnable for our hardware.
+The network takes sinusoidal embeddings of the progress coordinate as input and directly outputs the -dimensional molecular configuration, where is the number of atoms in the system.
+The standard MLP (TransitionMLP) uses the same interpolation strategy and endpoint constraints as AdaPath but processes all atomic coordinates through shared dense layers rather than using per-atom embeddings with parameter sharing.
+All other training parameters remain identical to the base configuration described in Section F.1, with both architectures using the same learning rate of 1e-2 with the Muon optimizer for the two smaller systems and the Nadamw optimizer for BPTI with the learning rate scheduling as outlined in G.
+| Architecture | Parameters | Max Energy | MinMax Energy | Mean Energy |
+| Standard MLP | 5e4 | |||
+| AdaPath | 2e6 |
+| Architecture | Parameters | Max Energy | MinMax Energy | Mean Energy |
+| Standard MLP | 2e6 | |||
+| AdaPath | 2e6 |
+| Architecture | Parameters | Max Energy | MinMax Energy | Mean Energy |
+| Standard MLP | 1.4e9 | |||
+| AdaPath | 3.5e6 |
+The results demonstrate AdaPath’s superior performance and scalability across system sizes.
+The AdaPath’s parameter count is dominated by the progress coordinate MLP that outputs the adaLn parameters.
+This produces a baseline parameter count that puts the AdaPath architecture above the MLP for small systems, but for larger ones, this gap inverts sharply.
+On the alanine dipeptide system, AdaPath achieves significantly lower and more consistent maximum energies (-86.6 vs -68.1 kJ/mol for minimum maximum energy) with substantially reduced variance.
+The performance gap becomes dramatic on the larger AIB9 system, where the standard MLP achieves a best maximum energy of 2.99e4 kJ/mol compared to AdaPath’s -644 kJ/mol, representing a 30,000+ kJ/mol improvement in path quality.
+F.3 Optimizer Comparison: Path Continuity Analysis
+To validate our claim that the Muon optimizer maintains better path continuity compared to Adam for small systems, we conduct a comparative analysis using path velocity metrics over training iterations.
+We train identical AdaPath networks on the AIB9 system using both optimizers with learning rates of 1e-2 for Muon and 1e-4 for Adam, respectively.
+All other training parameters remain identical to the base configuration.
+We track the maximum path velocity over training iterations, computed as , where high velocities indicate rapid changes in molecular configuration that can lead to discontinuous ”teleportation” between states rather than smooth transitions.
+| Optimizer | Max Energy | Mean Energy |
+| Adam | ||
+| Muon |
+The Adam-trained system exhibits extremely high path velocities, indicating that it rapidly switches from one state to another without passing through the transition region.
+This is corroborated by the low mean but high max energies.
+F.4 Energy Transformation Analysis
+We evaluate the impact of our logarithmic energy transformation on training stability by comparing performance with and without the log transformation applied to the potential energy in the loss function.
+The logarithmic transformation is applied as where ensures positivity, while the baseline uses the direct energy .
+All other training parameters remain identical to the base configuration.
+| Energy Transform | Max Energy | MinMax Energy | Mean Energy |
+| Direct Energy | |||
+| Log Transform |
+The logarithmic transformation achieves significantly lower energies with reduced variance, demonstrating improved training stability and path quality over direct energy optimization.
+F.4.1 Reference Trajectory Generation
+To generate reference conformational data, we performed an extended molecular dynamics simulation using OpenMM with the following parameters:
+Force field: AMBER ff15ipq-m
+Temperature: 500 K (elevated to enhance sampling of conformational transitions)
+Integrator: Langevin Middle Integrator with friction coefficient of 1 ps-1
+Time step: 1 femtosecond
+Non-bonded treatment: No cutoff (NoCutoff method)
+Constraints: Hydrogen bonds constrained using the HBonds method
+Total simulation time: 900 nanoseconds
+Output frequency: Coordinates saved every 1000 steps (1 ps intervals)
+The simulation protocol consisted of initial energy minimization followed by velocity assignment at the target temperature, a brief 10 ps equilibration period, and then the production run.
+The elevated temperature of 500 K was chosen to accelerate conformational sampling and ensure adequate transitions between metastable states during the simulation timeframe.
+F.4.2 State Definition and Analysis
+To characterize the conformational landscape, we focused on the backbone dihedral angles of the central residues.
+Specifically, we computed the and dihedral angles for residues at indices [56, 58, 60, 69] and [58, 60, 69, 71], respectively, corresponding to the central amino acids of the peptide chain.
+Two metastable states were defined using ellipsoidal boundaries in the - dihedral space:
+These elliptical regions capture the distinct conformational basins observed in the free energy landscape projected onto the central residue dihedral angles.
+F.4.3 Neural Network Training Setup
+For the neural network-based MEP optimization, we employed a different force field setup optimized for differentiable molecular dynamics:
+Force field: AMBER ff15ipq-m for protein mimetics, implemented through DMFF
+Non-bonded treatment: No cutoff method for consistency with the reference simulation
+Simulation box: 11.0 nm cubic box
+Neighbor list cutoff: 5.0 nm
+F.4.4 Start- and End-state Setup
+Start and end configurations for MEP optimization were randomly sampled from the reference trajectory frames that satisfied the respective state definitions.
+The Kabsch algorithm was employed to optimally align the end configuration to the start configuration, removing translational and rotational degrees of freedom.
+This alignment procedure ensures that the neural network focuses on learning the internal conformational changes rather than rigid-body motions, leading to more efficient and physically meaningful transition path discovery.
+Appendix G BPTI System Preparation
 As reference conformations, we used the structures from the D.E.
@@
 The initial and final states for our MEP optimization were selected as the second snapshot (index 1) and the second-to-last snapshot (index -2) from the provided reference structures, as they were the most separated in the CVs determined in the original study, named the backbone RMSD and the disulfide torsion angle.
-B.2.2 Explicit Solvent Shell Construction
-For the BPTI system, we developed a sophisticated approach to handle explicit water molecules efficiently:
-Initial water box generation: Both start and end configurations were solvated with TIP3P water molecules using a padding distance of 2.0 nm and energy minimized for 15,000 steps.
-Water shell extraction: Rather than using the entire water box, we extracted spherical shells of water molecules within a 1.4 nm radius of any protein atom for both the start and end configurations.
-This significantly reduced the system size while maintaining the crucial first hydration shell.
-Radius optimization: We employed a binary search algorithm to determine optimal shell radii that would capture an identical number of water molecules in both start and end configurations, ensuring consistent system sizes throughout the transition.
-Water molecule matching: We applied the Hungarian algorithm to optimally pair water molecules between the start and end states based on spatial proximity, minimizing the overall displacement of water molecules during the transition, which ensures that the intial guess for the MEP does not have very large energies stemming from overlapping molecules.
-B.2.3 Hybrid Neural Network Architecture
-For the BPTI system with explicit solvent, we developed a specialized hybrid transition model to handle the larger system size efficiently:
-Protein representation: Neural network with 6 layers, a width factor of 4, and sinusoidal time embedding for the protein atoms (58 residues, 892 atoms)
-Water representation: Parameterized spline interpolation between matched water molecules
-Control points: 7 control points for the spline interpolation
-Water weight: 0.1 (relative weight of water molecule contributions to the loss)
-Clamp threshold: (minimum/maximum energy clipping threshold)
-B.2.4 Training Parameters
-The BPTI MEP was optimized using the following training parameters:
-Optimizer: Adam
+G.1 Explicit Solvent System Preparation
+For the BPTI system, we developed an approach to handle explicit water molecules and ions using periodic boundary conditions:
+System solvation: Both start and end protein configurations were placed in 3.5 nm cubic periodic boxes and solvated with TIP3P water molecules using OpenMM’s Modeller class.
+The system included Na+ and Cl- ions at 0.04 M ionic strength to approximate physiological conditions.
+Energy minimization: To preserve the reference protein conformations from the D.E.
+Shaw simulation while allowing solvent relaxation, all protein heavy atoms were fixed by setting their masses to zero during energy minimization.
+This completely freezes the protein structure while allowing water molecules and ions to relax.
+Molecule matching: Since the endpoint configurations were solvated independently, we equalized atom counts by randomly removing excess molecules, then used the Hungarian algorithm to optimally pair water molecules and ions between configurations based on center-of-mass positions.
+This matching minimizes solvent displacement during transitions.
+The resulting system contained approximately 3,500 atoms total in a periodic cubic box, maintaining proper solvation while using standard periodic boundary conditions compatible with conventional MD force fields.
+G.2 Training Parameters
+The BPTI MEP optimization employed a two-stage training strategy due to spiking loss late during training.
+This is likely due to sharp minima in large molecular configurations, including water, where small changes in atomic coordinates can produce drastic energy spikes.
+This necessitated transitioning to more conservative learning parameters in Stage 2 to maintain training stability.
+Initial Training, iterations 0-6,000:
 Learning rate:
+Gradient clipping threshold:
+Fine-tuning, iterations 6,000-12,000:
+Learning rate:
+Gradient clipping threshold:
+Architecture and General Training Parameters:
+MLP Expansion Factor: 2
+Layers: 5
+Hidden dimension: 64
+Progress embedding dimension: 1024
+Conditioning depth: 3 layers
+Conditioning hidden size: 1024
+Optimizer: NAdam with weight decay (nadamw)
+Weight decay:
+Beta parameters: , ,
 Batch size: 1
 Frames per batch: 16
-Total iterations: 30,000
-All other parameters are identical to the base settings listed in the ablations section B.1.1.
-B.2.5 Handling High-Energy Spikes During Training
-Physical force fields present unique optimization challenges due to occasional extreme energy spikes when atoms approach one another too closely.
-To address this, we implemented adaptive gradient clipping in our training procedure:
-where is a clipping threshold (set to in our experiments).
-We specifically use adaptive gradient clipping rather than a fixed threshold because our hybrid architecture contains two fundamentally different parameter types: neural network weights for protein coordinates and spline control points for water molecules.
-The adaptive nature of the clipping automatically adjusts to the appropriate scale for each parameter type.
-Additionally, we apply different scaling factors to these parameter groups:
-with for water-related parameters.
-This differential treatment further balances the influence of each component during optimization.
-While this approach does not completely eliminate training instabilities, we find it sufficient for most protein systems when combined with careful initialization and learning rate scheduling.
-B.2.6 Force Field Parameters
+Total iterations: 12,000
+Evaluation frequency: Every 1000 iterations
+G.3 Force Field Parameters
 We used the following force field parameters to match the D.E. Shaw simulation conditions:
@@
 Water force field: TIP3P (tip3p.xml)
-B.2.7 Computational Performance
-For the BPTI system with explicit solvent (approximately 8,300 atoms total):
-Training time: Approximately 7 to 15 minutes on a single NVIDIA A6000 GPU
-Force field evaluations: Approximately 480,000 evaluations (30,000 iterations x 16 frames per batch)
+Appendix H Training Stability
+Atomistic force fields present unique optimization challenges due to the presence of Lennard-Jones interactions, which are typically modeled using the 12-6 potential:
+where is the distance between atoms and , is the well depth, and is the collision diameter.
+The repulsive term creates extremely steep energy barriers when atoms approach each other too closely, leading to numerical instabilities during neural network training.
+As , the potential energy diverges to infinity, causing gradient magnitudes that can destabilize the optimization process.
+To mitigate these issues, we employ a two-pronged approach. First, we apply a logarithmic transformation to the potential energy after shifting it to ensure positivity:
+Where is a positive constant to ensure the argument of the logarithm remains positive.
+This transformation preserves the location of energy minima while significantly compressing the magnitude of high-energy spikes.
+The monotonic nature of the logarithm ensures that , preserving the optimization objective.
+However, the model now places different weights on different parts of the path than before.
+In practice, we find that this drastically improves training stability.
+Second, we implement gradient norm-based gradient clipping to handle any remaining numerical instabilities.
+This combination of logarithmic energy transformation and gradient clipping provides robust training stability while maintaining the physical validity of the discovered minimum-energy paths.
+H.1 Logarithmic Energy Transformation Perturbation Analysis
+Consider the true minimum energy path for satisfying:
+Boundary conditions: and
+Orthogonality condition:
+where is the unit tangent vector along the path.
+Consider a perturbed path where and (fixed endpoints). At each point, decompose the perturbation:
+where and .
+Under perturbation, the energy at each point becomes:
+where is the Hessian of , and we examine the perpendicular components of the perturbation and how they interact with the Hessian, as these are the components of interest in this context.
+Since at the MEP:
+The change in the discretized original loss:
+Substituting the energy expansion:
+For the log-transformed loss:
+Using for small :
+Both losses exhibit:
+1.
+No first-order response to perpendicular perturbations: The terms appear only at second order
+2.
+Positive-definite second-order response: Since the MEP follows energy valleys,
+3.
+Restoring forces: Both and for any
+The difference lies only in the weighting factor , which scales the magnitude but not the direction of the restoring force. Therefore, both optimizations converge to the same geometric path.
+H.1.1 Parameterization and Path Stretching
+While both losses converge to the same MEP geometry, they differ in how they parameterize the path. The parallel component of the perturbation affects the path’s parameterization speed .
+The gradient with respect to shifting a sample point along the path:
+This creates forces that push sample points away from high-energy regions, potentially leading to extreme stretching where at transition states.
+The corresponding gradient for the log loss:
+The weighting factor provides implicit regularization:
+At transition states (high ): The factor is small, reducing the incentive to stretch
+At stable states (low ): The factor is larger but bounded
+Appendix I Computational Performance
+For the BPTI system with explicit solvent (approximately 3,500 atoms total):
+Training time: Approximately 15 minutes on a single NVIDIA A6000 GPU
+Force field evaluations: Approximately 192,000 evaluations (30,000 iterations x 16 frames per batch)
 Trajectory generation: 64 frames for the final MEP visualization and analysis
@@
 Shaw Research simulation on their specialized Anton supercomputer, which required approximately 1.3 million node-hours to generate the 1.03 millisecond trajectory capturing the same conformational changes.
-B.2.8 Computational Resources
+Appendix J Computational Resources
 All experiments reported in this paper were conducted on a single NVIDIA A6000 GPU with 48GB of memory.